### v3

# skipabove=4pt, skipbelow=4pt, leftmargin=0pt, rightmargin=0pt, innertopmargin=3pt, innerbottommargin=3pt, innerleftmargin=3pt, innerrightmargin=3pt, frametitleaboveskip=3pt, frametitlebelowskip=3pt, frametitlerule=true, splittopskip=2

Abstract
Generalized principal-agent problems, including Stackelberg games, contract design, and Bayesian persuasion, are a class of economic problems where an agent best responds to a principal’s committed strategy. We study repeated generalized principal-agent problems under the assumption that the principal does not have commitment power and the agent uses algorithms to learn to respond to the principal. We reduce this problem to a one-shot generalized principal-agent problem with an approximately-best-responding agent. Using this reduction, we show that: (1) if the agent uses contextual no-regret learning algorithms, then the principal can guarantee a utility that is at least the principal’s optimal utility in the classic non-learning model minus the square root of the agent’s regret; (2) if the agent uses contextual no-swap-regret learning algorithms, then the principal cannot obtain any utility more than the optimal utility in the non-learning model plus the agent’s swap regret. But (3) if the agent uses mean-based learning algorithms (which can be no-regret but not no-swap-regret), then the principal can do significantly better than the non-learning model. These general results not only refine previous results in Stackelberg games and contract design with learning agents but also lead to new results for Bayesian persuasion with a learning agent.
1 Introduction
Economic models involving principal-agent interactions, including auction design, contract design, and Bayesian persuasion, often assume that the agent is able to best respond to the strategy committed by the principal. For example, in Bayesian persuasion, the agent/receiver needs to compute the posterior belief about the state of the world after receiving the principal/sender’s information and then take an optimal action based on the posterior belief; this requires the receiver accurately knowing the prior of the state as well as the signaling scheme of the sender. The best-responding behavior, requiring strong rationality assumptions on the agent, is often observed to be violated in practice (Camerer, 1998; Benjamin, 2019).
In this work, using Bayesian persuasion as the main example, we study general principal-agent problems under an alternative behavioral model for the agent: learning. The use of learning as a behavioral model dates back to early economic literature on learning in games (Brown, 1951; Fudenberg and Levine, 1998) and has been actively studied recently in the computer science community (Nekipelov et al., 2015; Braverman et al., 2018; Deng et al., 2019; Mansour et al., 2022; Cai et al., 2023; Rubinstein and Zhao, 2024; Guruganesh et al., 2024). A learning agent no longer has perfect knowledge of the parameter of the game or the principle’s strategy. Instead of best responding, which is no longer possible or well-defined, the agent chooses his action based on past interactions with the principal. We focus on no-regret learning, which requires the agent to not feel a large regret due to not taking the optimal action in the end of repeated interactions with the principal. This is a mild requirement satisfied by many natural learning algorithms (e.g., $\varepsilon$-greedy, MWU, UCB, Exp3) and can reasonably serve as a possible behavioral assumption for real-world agents.
Our contributions
We first define a model of generalized principal-agent problem (Myerson, 1982; Gan et al., 2024) with a learning agent. This includes Stackelberg games, contract design, and Bayesian persuasion with a learning agent as special cases.
Then, we characterize the principal’s optimal strategy and payoff in generalized principal-agent problems with a learning agent. Our first result is that, against a no-regret learning agent, the principal can always guarantee a payoff that is arbitrarily close to the principal’s optimal payoff in the original principal-agent problem where the agent best responds. The difference between the principal’s payoff in the learning model and the best-response model is at most $O\big{(}\sqrt{\tfrac{\mathrm{Reg}(T)}{T}}\big{)}$ where $\mathrm{Reg}(T)$ is the regret of the agent and $T$ is the time horizon:
Theorem 1.1 (Informal).
Playing against a no-regret learning agent with regret $\mathrm{Reg}(T)$ in $T$ periods, the principal can always obtain an average payoff of at least $U^{*}-O\big{(}\sqrt{\tfrac{\mathrm{Reg}(T)}{T}}\big{)}$, where $U^{*}$ is the principal’s optimal payoff in the principal-agent problem with best response. To achieve this, the principal does not need to know the exact learning algorithm of the agent (knowing the regret bound $\mathrm{Reg}(T)$ is sufficient) and the principal can simply use a fixed strategy across all periods.
Our second result shows that, if the agent does a stronger version of no-regret learning, no-swap-regret learning (Hart and Mas-Colell, 2000; Blum and Mansour, 2007), then the principal can never obtain any payoff that is significantly higher than the optimal payoff in the best-response model.
Theorem 1.2 (Informal).
Playing against a no-swap-regret learning agent with swap-regret $\mathrm{SReg}(T)$ in $T$ periods, the principal cannot obtain any payoff larger than $U^{*}+O\big{(}\tfrac{\mathrm{SReg}(T)}{T}\big{)}$. This holds even if the principal knows the learning algorithm of the agent and uses time-varying strategies.
Finally, we show that, if the agent uses some learning algorithm satisfying no-regret but not no-swap-regret, then there exist instances where the principal can do significantly better than the best-response optimal objective $U^{*}$. We show this in the Bayesian persuasion problem specifically:
Theorem 1.3 (Informal).
There exists a Bayesian persuasion instance where, against a no-regret but not no-swap-regret learning agent, the principal can do significantly better than $U^{*}$.
Our general results above, when applied to Stackelberg games and contract design, refine the results of Deng et al. (2019) and Guruganesh et al. (2024) by bounding the difference between $U^{*}$ and the utility that can be obtained by the principal against a learning agent. When applied to Bayesian persuasion, we refine the result of Jain and Perchet (2023) (who study no-regret and mean-based learning agents in Bayesian persuasion concurrently and independently of us) and additionally show that the agent’s no-swap-regret learning can cap the principal’s utility. The conclusion that no-swap-regret learning can cap the principal’s utility was known in Stackelberg games and contract design. We show that it in fact holds for the much larger class of all generalized principal-agent problems with complete information.
Reduction from learning to approximate best response
The main intuition behind our first two results is: principal-agent problems with a no-regret learning agent is closely related principal-agent problems with an agent who approximately best respond to the principal’s strategy. The no-regret learning behavior of the agent can be seen as a form of approximate best response, where the agent makes sub-optimal decisions with the sub-optimality measured by the regret. When the sub-optimality/regret is small, the principal-agent problem with approximate best response returns to the problem with exact best response. This explains why the principal against a no-regret learning agent can obtain a payoff that is close to the payoff with a best-responding agent, and cannot obtain any payoff better than that.
However, formalizing the exact connection between learning and approximate best response is tricky. In particular, since no-regret (no-swap-regret) learning algorithms are randomized, they correspond to approximately-best-responding agents who use randomized strategies. We find that, against an agent using randomized $\delta$-approximately-best-responding strategies, the principal can and can only obtain payoff in $[U^{*}-O(\sqrt{\delta}),U^{*}+O(\delta)]$. If the agent only uses deterministic $\delta$-approximately-best-responding strategies, then the range of principal payoff is $[U^{*}-O(\delta),U^{*}+O(\delta)]$. Here, we see an interesting asymmetry between the lower bound (with a squared-root dependency on $\delta$) and the upper bound (with a linear dependency on $\delta$) with randomized strategies, but not with deterministic strategies.
Structure of the paper
We define the model of generalized principal-agent problems with a learning agent in Section 2. Since Bayesian persuasion is the main motivation of our work, we also present the specific model of persuasion with a learning agent in Section 2.3. Interestingly, Bayesian persuasion and cheap talk are equivalent under our learning agent model. We develop our main results in Section 3 and Section 4, by first reducing the generalized principal-agent problem with a learning agent to the problem with approximate best response, and then characterizing the problem with approximate best response. Then, Section 5 applies our general results to three specific principal-agent problems: Bayesian persuasion, contract design, and Stackelberg games. Section 6 offers additional discussions.
1.1 Related Works
The literature on information design (Bayesian persuasion) has investigated various relaxations of the strong rationality assumptions in the classic models. For the sender, known prior (Dworczak and Pavan, 2020; Camara et al., 2020; Kosterina, 2022; Zu et al., 2021; Ziegler, 2020; Wu et al., 2022) and known utility (Babichenko et al., 2021; Castiglioni et al., 2020; Feng et al., 2022) are relaxed. For the receiver, the receiver may make mistakes in Bayesian updates (de Clippel and Zhang, 2022), be risk-conscious (Anunrojwong et al., 2023), do quantal response (Feng et al., 2024) or approximate best response (Yang and Zhang, 2024). We focus on a learning receiver and show that it can be reduced to approximate best response. While Yang and Zhang (2024) show that Bayesian persuasion with approximate best response is computationally hard as the instance size increases, our results imply that this problem is not hard with a fixed instance and a diminishing degree of approximate best response.
Learning agents have been studied in principal-agent problems like auctions (Braverman et al., 2018; Cai et al., 2023; Rubinstein and Zhao, 2024; Kumar et al., 2024), bimatrix Stackelberg games (Deng et al., 2019; Mansour et al., 2022; Arunachaleswaran et al., 2024), contract design (Guruganesh et al., 2024), and Bayesian persuasion (Jain and Perchet, 2023). These problems belong to the class of generalized principal-agent problems (Myerson, 1982; Gan et al., 2024). We thus propose a general framework of generalized principal-agent problem with a learning agent, which encompasses several previous models, refines known results, and provides new results.
Camara et al. (2020) also propose a general framework of principal-agent problems with learning players, but has two key differences with ours: (1) They drop the common prior assumption while we still keep it. This assumption allows us to compare the principal’s utility in the learning model with the classic model with common prior. (2) Their principal has commitment power, which is reasonable in, e.g., auction design, but less realistic in information design where the principal’s strategy is a signaling scheme. Our principal does not commit.
Deng et al. (2019) show that the follower’s no-swap-regret learning can cap the leader’s utility in Stackelberg games. We find that this conclusion holds for all generalized principal-agent problems with complete information. Nevertheless, this conclusion does not hold when the agent has private information, as shown by Mansour et al. (2022)’s work on Bayesian Stackelberg games. We believe that generalized principal-agent problems with private information with learning players is a viable direction for future research.
2 Model: Generalized Principal-Agent Problem with a Learning Agent
This section defines our model, generalized principal-agent problem with a learning agent. This model includes Stackelberg games, contract design, and Bayesian persuasion with learning agents.
2.1 Generalized Principal-Agent Problem
Generalized principal-agent problem, originally proposed by (Myerson, 1982), is a general model that includes auction design, contract design, and Stackelberg games. Gan et al. (2024) further generalize it to include Bayesian persuasion. We take the complete-information variant of (Gan et al., 2024)’s model.111 The models of Myerson (1982) and Gan et al. (2024) allow the agent to have private information. Our work studies a complete-information model. Our conclusions do not extend to the private-information case.
There are two players in a generalized principal-agent problem: a principal and an agent. The principal has a convex, compact decision space $\mathcal{X}$ and the agent has a finite action set $A$. The principal and the agent have utility functions $u,v:\mathcal{X}\times A\to{\mathbb{R}}$. Assume that $u(x,a)$, $v(x,a)$ are linear in $x\in\mathcal{X}$, which is satisfied by all the examples of generalized principal-agent problems we will consider (Bayesian persuasion, Stackelberg games, contract design). There is a signal/message set $S$. Signals are usually interpreted as action recommendations, where $S=A$, but we allow any general finite signal set. A strategy of the principal is a distribution $\pi\in\Delta(\mathcal{X}\times S)$ over pairs of decision and signal. When the utility functions $u,v$ are linear, it is without loss of generality to assume that the principal does not randomize over multiple decisions for one signal, namely, the principal chooses a distribution over signals and a unique decision $x_{s}$ associated with each signal $s\in S$ (Gan et al., 2024). So, we can write a principal strategy as $\pi=\{(\pi_{s},x_{s})\}_{s\in S}$ where $\pi_{s}\geq 0$ is the probability of signal $s\in S$, $\sum_{s\in S}\pi_{s}=1$, and $x_{s}\in\mathcal{X}$. There are two variants of generalized principal-agent problems:
  • Unconstrained (Myerson, 1982): there is no restriction on the principal’s strategy $\pi$.
  • Constrained (Gan et al., 2024): the principal’s strategy $\pi$ has to satisfy the constraint:
    $\sum_{s\in S}\pi_{s}x_{s}\in\mathcal{C}$ (1)
    where $\mathcal{C}\subseteq\mathcal{X}$ is some convex set.
Unconstrained generalized principal-agent problems include contract design and Stackelberg games (see Section 5 for details). Constrained generalized principal-agent problems further include Bayesian persuasion (discussed later).
In a one-shot generalized principal-agent problem where the principal has commitment power, the principal first commits to a strategy $\pi=\{(\pi_{s},x_{s})\}_{s\in S}$, then nature draws a signal $s\in S$ according to the distribution $\{\pi_{s}\}_{s\in S}$ and sends $s$ to the agent (note: due to the commitment assumption, this is equivalent to revealing the pair $(s,x_{s})$ to the agent), then the agent takes an action $a_{s}\in\operatorname*{arg\,max}_{a\in A}v(x_{s},a)$ to maximize its utility, and the principal obtains utility $u(x_{s},a_{s})$. The principal aims to maximize its expected utility $\mathbb{E}_{s\sim\pi}[u(x_{s},a_{s})]$ by choosing the strategy $\pi$.
2.2 Learning Agent
Now we define the model of generalized principal-agent problem with a learning agent. The game is repeated for $T$ rounds. The principal has no commitment power. The agent does not know the strategy $\pi^{t}$ or the decision $x^{t}$ of the principal at each round. Instead, the agent uses some adaptive algorithm to learn which action to take in response to each possible signal. {mdframed}[frametitle=Generalized Principal-Agent Problem with a Learning Agent] In each round $t=1,\ldots,T$:
  • (1)
    The principal chooses a strategy $\pi^{t}=\{(\pi_{s}^{t},x_{s}^{t})\}_{s\in S}$, which is a distribution over signals $S$ and a decision $x_{s}^{t}\in\mathcal{X}$ associated with each signal. The agent does not know $\pi^{t}$.
  • (2)
    Nature draws signal $s^{t}\sim\pi^{t}$ and reveals it. The principal makes decision $x^{t}=x_{s^{t}}^{t}$.
  • (3)
    Based on $s^{t}$ and history (and not knowing $x^{t}$), the agent (randomly) chooses an action $a^{t}\in A$, using some learning algorithm (described in details later).
  • (4)
    The principal and the agent obtain utility $u^{t}=u(x^{t},a^{t})$ and $v^{t}=v(x^{t},a^{t})$. The agent observes some feedback (which can be different from $v^{t}$).
With some knowledge of the agent’s learning algorithm, the principal aims to maximize its expected average utility $\frac{1}{T}\mathbb{E}\big{[}\sum_{t=1}^{T}u(x^{t},a^{t})\big{]}$.
Agent’s learning algorithm: no-regret and no-swap-regret
The agent’s learning problem can be regarded as a contextual multi-armed bandit problem (Tyler Lu et al., 2010) where $A$ is the set of arms, and the signal $s^{t}\in S$ serves as a context that affects the utility of each arm $a\in A$. The agent picks an arm to pull based on the current context $s^{t}$ and the historical information about each arm under different contexts, adjusting its strategy over time based on the feedback collected after each round.
What feedback can the agent observe after each round? One may assume that the agent sees the principal’s decision $x^{t}$ after each round (this is call full-information feedback in the multi-armed bandit literature), or the utility $v^{t}=v(x^{t},a^{t})$ obtained in that round (this is called bandit feedback), or even just some unbiased estimate of $v(x^{t},a^{t})$. We do not make any specific assumptions on the feedback observed by the agent. All we need is that the feedback is sufficient for the agent to achieve contextual no-regret or contextual no-swap-regret, which are defined below.
Definition 2.1.
The agent’s learning algorithm satisfies:
  • contextual no-regret if: there exists a function $\mathrm{CReg}(T)=o(T)$ such that, for any deviation function $d:S\to A$, the regret of the agent not deviating according to $d$ is at most $\mathrm{CReg}(T)$:
    $\mathbb{E}\Big{[}\sum_{t=1}^{T}\Big{(}v(x^{t},d(s^{t}))-v(x^{t},a^{t})\Big{)}\Big{]}\,\leq\,\mathrm{CReg}(T)=o(T),\quad\quad\forall d:S\to A.$ (2)
  • contextual no-swap-regret if: there exists a function $\mathrm{CSReg}(T)=o(T)$ such that, for any deviation function $d:S\times A\to A$, the regret of the receiver not deviating according to $d$ is at most $\mathrm{CSReg}(T)$:
    $\mathbb{E}\Big{[}\sum_{t=1}^{T}\Big{(}v(x^{t},d(s^{t},a^{t}))-v(x^{t},a^{t})\Big{)}\Big{]}\,\leq\,\mathrm{CSReg}(T)=o(T),\quad\quad\forall d:S\times A\to A.$ (3)
We call $\mathrm{CReg}(T)$ and $\mathrm{CSReg}(T)$ the contextual regret and contextual swap-regret of the agent.
Clearly, contextual no-regret is implied by contextual no-swap-regret because the latter has a larger set of deviation functions. Contextual no-(swap-)regret algorithms can be constructed easily by running an ordinary no-(swap-)regret algorithm for each context independently. Since algorithms with $O(\sqrt{T})$ (swap-)regret with bandit feedback are known to exist (Audibert and Bubeck, 2010; Ito, 2020), they can be converted into algorithms with $O(\sqrt{|S|T})$ contextual (swap-)regret. This is formalized by the following proposition, with proof in Appendix A.
Proposition 2.1.
There exist learning algorithms with contextual regret $\mathrm{CReg}(T)=O(\sqrt{|A||S|T})$ and contextual swap-regret $\mathrm{CSReg}(T)=O(|A|\sqrt{|S|T})$. They can be constructed by running an ordinary no-(swap-)regret multi-armed bandit algorithm for each context independently.
2.3 Special Case: Bayesian Persuasion with a Learning Agent
We will show in Section 5 that Stackelberg games and contract design are special cases of generalized principal-agent problems. Here, we explain why Bayesian persuasion is a special case of constrained generalized principal-agent problems. We will also show that Bayesian persuasion is in fact equivalent to cheap talk (Crawford and Sobel, 1982) under our learning agent model.
Bayesian persuasion as a generalized principal-agent problem
There are two players in Bayesian persuasion (Kamenica and Gentzkow, 2011): a sender (principal) and a receiver (agent). There are a finite set $\Omega$ of states of the world, a signal set $S$, an action set $A$, a prior distribution $\mu_{0}\in\Delta(\Omega)$ over the states, a utility function $u:\Omega\times A\to{\mathbb{R}}$ for the sender, and a utility function $v:\Omega\times A\to{\mathbb{R}}$ for the receiver. When the state is $\omega\in\Omega$ and the receiver takes action $a\in A$, the sender and the receiver obtain utility $u(\omega,a)$, $v(\omega,a)$, respectively. Both players know $\mu_{0}$, but only the sender has access to the realized state $\omega\sim\mu_{0}$. The sender commits to some signaling scheme $\pi:\Omega\to\Delta(S)$, mapping any state to a probability distribution over signals, to partially reveal information about the state $w$ to the receiver. In the classic model, after receiving a signal $s\in S$, the receiver will form the posterior belief $\mu_{s}\in\Delta(\Omega)$ about the state:
$\mu_{s}(\omega)=\frac{\mu_{0}(\omega)\pi(s|\omega)}{\pi_{s}},\text{ where }\pi_{s}=\sum_{\omega\in\Omega}\mu_{0}(\omega)\pi(s|\omega)\text{is the total probability that signal $s$ is sent.}$.
and take an optimal action with respect to $\mu_{s}$, $a_{s}\in\operatorname*{arg\,max}_{a\in A}\sum_{\omega\in\Omega}\mu_{s}(\omega)v(\omega,a)$. The sender aims to find a signaling scheme to maximizde its expected utility $\mathbb{E}[u(w,a_{s})]$.
It is well-known (Kamenica and Gentzkow, 2011) that a signaling scheme $\pi:\Omega\to\Delta(S)$ in the Bayesian persuasion problem decomposes the prior $\mu_{0}$ into a distribution over posteriors:
$\sum_{s\in S}\pi_{s}\mu_{s}=\mu_{0}\in\{\mu_{0}\}=:\mathcal{C},\quad\sum_{s\in S}\pi_{s}=1.$ (4)
Equation (4) is called the Bayes plausibility condition. Conversely, any distribution over posteriors $\{(p_{s},\mu_{s})\}_{s\in S}$ satisfying Bayes plausibility $\sum_{s\in S}p_{s}\mu_{s}=\mu_{0}$ can be converted into a signaling scheme that sends signal $s$ with probability $p_{s}$. Thus, we can use a distribution over posteriors $\{(\pi_{s},\mu_{s})\}_{s\in S}$ satisfying Bayes plausibility to represent a signaling scheme. Then, let’s equate the posterior belief $\mu_{s}$ in Bayesian persuasion to the principal’s decision $x_{s}$ in the generalized principal-agent problem, so the principal/sender’s decision space is $\mathcal{X}=\Delta(\Omega)$. The Bayes plausibility condition (4) becomes the constraint in the constrained generalized principal-agent problem. When the agent/receiver takes action $a$, the principal/sender’s (expected) utility under decision/posterior $x_{s}=\mu_{s}$ is:
$u(x_{s},a)=\mathbb{E}_{\omega\sim\mu_{s}}u(\omega,a)=\sum_{\omega\in\Omega}\mu_{s}(\omega)u(\omega,a).$
Suppose the agent takes action $a_{s}$ given signal $s\in S$. Then we see that the sender’s utility of using signaling scheme $\pi$ in Bayesian persuasion (left) is equal to the principal’s utility of using strategy $\pi$ in the generalized principal-agent problem (right):
$\sum_{\omega\in\Omega}\mu_{0}(\omega)\sum_{s\in S}\pi(s|\omega)u(\omega,a_{s})=\sum_{s\in S}\pi_{s}\sum_{\omega\in\Omega}\mu_{s}(\omega)u(\omega,a_{s})=\sum_{s\in S}\pi_{s}u(x_{s},a_{s})=\mathbb{E}_{s\sim\pi}[u(x_{s},a)].$
Similarly, the agent/receiver’s utilities in the two problems are equal. The utility functions $u(x,a)$, $v(x,a)$ are linear in the principal’s decision $x\in\mathcal{X}$, satisfying our assumption.
Persuasion and cheap talk with a learning agent
When specialized to Bayesian persuasion, the generalized principal-agent problem with a learning agent becomes the following:
{mdframed}
[frametitle=Persuasion with a Learning Receiver] In each round $t=1,\ldots,T$, the following events happen:
  • (1)
    The sender chooses a signaling scheme $\pi^{t}:\Omega\to\Delta(S)$, mapping each state to a distribution over signals. The receiver does not know $\pi^{t}$.
  • (2)
    A state of the world $\omega^{t}\sim\mu_{0}$ is drawn, observed by the sender but not the receiver. The sender then sends signal $s^{t}\sim\pi^{t}(\omega^{t})$ to the receiver.
  • (3)
    Upon receiving $s^{t}$, the receiver (randomly) chooses some action $a^{t}\in A$ to take, using some learning algorithm.
  • (4)
    The sender obtains utility $u^{t}=u(\omega^{t},a^{t})$ and the receiver obtains utility $v^{t}=v(\omega^{t},a^{t})$. [^2] The definition of utility here, $u(\omega^{t},a^{t}),v(\omega^{t},a^{t})$, is different from the definition in the previous model, which was the expected utility on decision/posterior $x^{t}$, $u(x^{t},a^{t}),v(x^{t},a^{t})$. Because we are only concerned with the sender’s expected utility and the receiver’s expected regret, this difference does not matter. The receiver observes some feedback.
As mentioned, the feedback observed by the receiver can take any form as long as the receiver is able to run contextual no-regret or no-swap-regret algorithms.
In the above model, the receiver does not need to know the prior $\mu_{0}$ or the sender’s signaling scheme. Since the receiver does not know the sender’s signaling scheme when making decisions, we can flip the order of decision making of the two players, obtaining an equivalent model: {mdframed}[frametitle=Cheap Talk with a Learning Agent] In each round $t=1,\ldots,T$, the following events happen:
  • (1’)
    The receiver first chooses, using some learning algorithm, a strategy $\rho^{t}:S\to\Delta(A)$ that maps each signal $s\in S$ to a probability distribution over actions $\rho^{t}(s)\in\Delta(A)$.
  • (2’)
    The sender chooses a signaling scheme $\pi^{t}:\Omega\to\Delta(S)$.
  • (3’)
    A state of the world $\omega^{t}\sim\mu_{0}$ is drawn. Signal $s^{t}\sim\pi^{t}(\omega^{t})$ is sent to the receiver, who then chooses an action $a^{t}$ according to the distribution $\rho^{t}(s^{t})$.
  • (4)
    The sender obtains utility $u^{t}=u(\omega^{t},a^{t})$ and the receiver obtains utility $v^{t}=v(\omega^{t},a^{t})$.
One can see from above that the sender does not need to commit to the signaling scheme. The receiver’s strategy $\rho^{t}$ is chosen before the sender chooses the signaling scheme, so the sender has an incentive to best respond to $\rho^{t}$. This corresponds to the well-known cheap talk model (Crawford and Sobel, 1982) where the sender does not have commitment power. So, the “persuasion with a learning receiver” model is equivalent to “cheap talk with a learning receiver”.
3 Reduction from Learning to Approximate Best Response
Our first main result is a reduction from the generalized principal-agent problem with a learning agent to the problem with an approximately-best-responding agent. We show that, if the agent uses contextual no-regret learning algorithms, then the principal can always obtain an average utility that is at least the “maximin” approximate-best-response objective $\underline{\mathrm{OBJ}}^{\mathcal{R}}\big{(}\mathrm{CReg}(T)/T\big{)}$ (to be defined below). On the other hand, if the agent does contextual no-swap-regret learning, then the principal cannot do better than the “maximax” approximate-best-response objective $\overline{\mathrm{OBJ}}^{\mathcal{R}}\big{(}\mathrm{CSReg}(T)/T\big{)}$. In addition, we show that the principal can do better than the “maximax” objective $\overline{\mathrm{OBJ}}^{\mathcal{R}}\big{(}\mathrm{CSReg}(T)/T\big{)}$ if the agent uses some natural learning algorithms that are no-regret but not no-swap-regret, in particular, mean-based algorithms (Braverman et al., 2018).
3.1 Generalized Principal-Agent Problem with Approximate Best Response
We first define a generalized principal-agent problem with an approximately-best-responding agent.
The classic generalized principal-agent problem (Section 2.1) assumes that, after receiving a signal $s\in S$ (and observing the principal’s decision $x_{s}\in\mathcal{X}$), the agent will take an action $\in\operatorname*{arg\,max}_{a\in A}v(x_{s},a)$ to maximize its utility with respect to $x_{s}$. This means that the agent uses a strategy $\rho^{*}$ that best responds to the principal’s strategy $\pi$:
$\rho^{*}(s)\in\operatorname*{arg\,max}_{a\in A}v(x_{s},a),~{}~{}\forall s\in S\quad\implies\quad\rho^{*}\in\operatorname*{arg\,max}_{\rho:S\to\Delta(A)}V(\pi,\rho).$ (5)
Here, $V(\pi,\rho)$ denotes the expected utility of the agent when the principal uses strategy $\pi$ and the agent uses (randomized) strategy $\rho:S\to\Delta(A)$:
$V(\pi,\rho)=\sum_{s\in S}\pi_{s}\sum_{a\in A}\rho(a|s)v(x_{s},a).$ (6)
Here, we allow the agent to approximately best respond. Let $\delta\geq 0$ be a parameter. We define two types of $\delta$-best-responding strategies for the agent: deterministic and randomized.
  • A deterministic strategy $\rho$: for each signal $s\in S$, the agent takes an action $a$ that is $\delta$-optimal for $x_{s}$. Denote this set of strategies by $\mathcal{D}_{\delta}(\pi)$:
    $\mathcal{D}_{\delta}(\pi)=\big{\{}\rho:S\to A\mid v(x_{s},\rho(s))\geq v(x_{s},a^{\prime})-\delta,~{}\forall a^{\prime}\in A\big{\}}.$ (7)
  • A randomized strategy $\rho$: for each signals $s$, the agent can take a randomized action. The expected utility of $\rho$ is at most $\delta$-worst than the best strategy $\rho^{*}$.
    $\mathcal{R}_{\delta}(\pi)=\big{\{}\rho:S\to\Delta(A)\mid V(\pi,\rho)\geq V(\pi,\rho^{*})-\delta\big{\}}.$ (8)
    Equivalently, $\mathcal{R}_{\delta}(\pi)=\big{\{}\rho:S\to\Delta(A)\mid V(\pi,\rho)\geq V(\pi,\rho^{\prime})-\delta,~{}\forall\rho^{\prime}:S\to A\big{\}}$.
Clearly, $\mathcal{D}_{\delta}(\pi)\subseteq\mathcal{R}_{\delta}(\pi)$.
Our model of approximately-best-responding agent includes, for example, two other models in the Bayesian persuasion literature that also relax the agent’s Bayesian rationality assumption: the quantal response model (proposed by McKelvey and Palfrey (1995) in normal-form games and studied by Feng et al. (2024) in Bayesian persuasion) and a model where the agent makes mistakes in Bayesian update (de Clippel and Zhang, 2022).
Example 3.1.
Assume that the receiver’s utility is in $[0,1]$. In Bayesian persuasion, the following receiver strategies are $\delta$-best-responding:
  • Quantal response: given signal $s\in S$, the agent chooses action $a\in A$ with probability $\frac{\exp(\lambda v(\mu_{s},a))}{\sum_{a^{\prime}\in A}\exp(\lambda v(\mu_{s},a^{\prime}))}$, with $\lambda>0$. This strategy belongs to $\mathcal{R}_{\delta}(\pi)$ with $\delta=\frac{1+\log(|A|\lambda)}{\lambda}$.
  • Inaccurate belief: given signal $s\in S$, the agent forms some posterior $\mu_{s}^{\prime}$ that is different yet close to the true posterior $\mu_{s}$ in total variation distance $d_{\mathrm{TV}}(\mu_{s}^{\prime},\mu_{s})\leq\varepsilon$. The agent picks an optimal action for $\mu_{s}^{\prime}$. This strategy belongs to $\mathcal{D}_{2\varepsilon}(\pi)$.
See Appendix B for a proof of this example.
Principal’s objectives
With an approximately-best-responding agent, we will study two types of objectives for the principal. The first type is the maximal utility that the principal can obtain if the agent approximately best responds in the worst way for the principal: for $X\in\{\mathcal{D},\mathcal{R}\}$, define
$\underline{\mathrm{OBJ}}^{X}(\delta)=\sup_{\pi}\min_{\rho\in X_{\delta}(\pi)}U(\pi,\rho),$ (9)
where $U(\pi,\rho)$ is the principal’s expected utility when the principal uses strategy $\pi$ and the agent uses strategy $\rho$:
$U(\pi,\rho)=\sum_{s\in S}\pi_{s}\sum_{a\in A}\rho(a|s)u(x_{s},a).$ (10)
We used “$\sup$” in (9) because the maximizer does not necessarily exist. $\underline{\mathrm{OBJ}}^{X}(\delta)$ is a “maximin” objective and can be regarded as the objective of a “robust generalized principal-agent problem”.
The second type of objectives is the maximal utility that the principal can obtain if the agent approximately best responds in the best way:
$\overline{\mathrm{OBJ}}^{X}(\delta)=\max_{\pi}\max_{\rho\in X_{\delta}(\pi)}U(\pi,\rho).$ (11)
This is a “maximax” objective that quantifies the maximal extent to which the principal can exploit the agent’s irrational behavior.
Clearly, $\underline{\mathrm{OBJ}}^{X}(\delta)\leq\underline{\mathrm{OBJ}}^{X}(0)\leq\overline{\mathrm{OBJ}}^{X}(0)\leq\overline{\mathrm{OBJ}}^{X}(\delta)$. And we note that $\overline{\mathrm{OBJ}}^{X}(0)=\overline{\mathrm{OBJ}}(0)$ is independent of $X$ and equal to the optimal principal utility in the classic generalized principal-agent problem, which we denote by $U^{*}$:
$U^{*}=\overline{\mathrm{OBJ}}(0)=\max_{\pi}\max_{\rho:\text{ best-response to $\pi$}}U(\pi,\rho).$ (12)
Finally, we note that, because $\mathcal{D}_{0}(\pi)\subseteq\mathcal{D}_{\delta}(\pi)\subseteq\mathcal{R}_{\delta}(\pi)$, the following series of inequalities hold:
$\underline{\mathrm{OBJ}}^{\mathcal{R}}(\delta)\leq\underline{\mathrm{OBJ}}^{\mathcal{D}}(\delta)\leq U^{*}\leq\overline{\mathrm{OBJ}}^{\mathcal{D}}(\delta)\leq\overline{\mathrm{OBJ}}^{\mathcal{R}}(\delta).$ (13)
3.2 Agent’s No-Regret Learning: Lower Bound on Principal’s Utility
Theorem 3.1.
Suppose the agent uses a contextual no-regret learning algorithm. By using some fixed strategy $\pi^{t}=\pi$ in all $T$ rounds, the principal can obtain an average utility $\frac{1}{T}\mathbb{E}\big{[}\sum_{t=1}^{T}u(x^{t},a^{t})\big{]}$ that is arbitrarily close to $\underline{\mathrm{OBJ}}^{\mathcal{R}}\big{(}\mathrm{CReg}(T)/T\big{)}$.
To prove Theorem 3.1, we first prove a lemma to characterize the agent’s regret and the principal’s expected utility. We define some notations. Let the principal use some fixed strategy $\pi^{t}=\pi$ and the agent use some learning algorithm. Let $p_{a|s}^{t}=\Pr[a^{t}=a\,|\,s^{t}=s]$ be the probability that the agent’s algorithm chooses action $a$ conditioning on signal $s$ being sent in round $t$. Let $\rho:S\to\Delta(A)$ be a randomized agent strategy that, given signal $s$, chooses each action $a\in A$ with probability
$\rho(a|s)=\tfrac{\sum_{t=1}^{T}p_{a|s}^{t}}{T}.$ (14)
Lemma 3.2.
When the principal uses a fixed strategy $\pi^{t}=\pi$ in all $T$ rounds, the regret of the agent not deviating according to $d:S\to A$ is equal to $\frac{1}{T}\mathbb{E}\big{[}\sum_{t=1}^{T}\big{(}v(x^{t},d(s^{t}))-v(x^{t},a^{t})\big{)}\big{]}=V(\pi,d)-V(\pi,\rho)$, and the average utility of the sender $\frac{1}{T}\mathbb{E}\big{[}\sum_{t=1}^{T}u(x^{t},a^{t})\big{]}$ is equal to $U(\pi,\rho)$.
Proof.
Since $\pi^{t}=\pi$ is fixed, we have $\pi_{s}^{t}=\pi_{s}$ and $x_{s}^{t}=x_{s}$, $\forall s\in S$. The regret of the receiver not deviating according to $d$ is:
$\frac{1}{T}\mathbb{E}\Big{[}\sum_{t=1}^{T}\Big{(}v(x^{t},d(s^{t}))-v(x^{t},a^{t})\Big{)}\Big{]}=\frac{1}{T}\sum_{t=1}^{T}\sum_{s\in S}\pi^{t}_{s}\sum_{a\in A}p_{a|s}^{t}\Big{(}v(x_{s}^{t},d(s))-v(x_{s}^{t},a)\Big{)}$
$~{}~{}=\sum_{s\in S}\pi_{s}\sum_{a\in A}\tfrac{\sum_{t=1}^{T}p_{a|s}^{t}}{T}\Big{(}v(x_{s},d(s))-v(x_{s},a)\Big{)}$
$~{}~{}=\sum_{s\in S}\pi_{s}v(x_{s},d(s))~{}-~{}\sum_{s\in S}\pi_{s}\sum_{a\in A}\rho(a|s)v(x_{s},a)~{}=~{}V(\pi,d)-V(\pi,\rho).$
Here, $d$ is interpreted as an agent strategy that deterministically takes action $d(s)$ for signal $s$.
By a similar derivation, we see that the principal’s expected utility is equal to
$\frac{1}{T}\mathbb{E}\Big{[}\sum_{t=1}^{T}u(x^{t},a^{t})\Big{]}=\sum_{s\in S}\pi_{s}\sum_{a\in A}\tfrac{\sum_{t=1}^{T}p_{a|s}^{t}}{T}u(x_{s},a)=U(\pi,\rho),$
which proves the lemma. ∎
Proof of Theorem 3.1.
By Lemma 3.2 and the no-regret condition that the agent’s regret
$\mathbb{E}\big{[}\sum_{t=1}^{T}\big{(}v(x^{t},d(s^{t}))-v(x^{t},a^{t})\big{)}\big{]}\leq\mathrm{CReg}(T)$, we have
$V(\pi,d)-V(\pi,\rho)\,=\,\frac{1}{T}\mathbb{E}\Big{[}\sum_{t=1}^{T}\Big{(}v(x^{t},d(s^{t}))-v(x^{t},a^{t})\Big{)}\Big{]}\,\leq\,\frac{\mathrm{CReg}(T)}{T},\quad\forall d:S\to A$
This means that the agent’s randomized strategy $\rho$ is a $\delta=\frac{\mathrm{CReg}(T)}{T}$-best-response to the principal’s fixed signaling scheme $\pi$:
$\rho\in\mathcal{R}_{\delta=\frac{\mathrm{CReg}(T)}{T}}(\pi).$
This holds for any $\pi$. In particular, if for any $\varepsilon>0$ the principal uses a signaling scheme $\pi^{\varepsilon}$ that obtains an objective that is $\varepsilon$-close to $\underline{\mathrm{OBJ}}^{\mathcal{R}}(\delta)=\sup_{\pi}\min_{\rho\in\mathcal{R}_{\delta}(\pi)}U(\pi,\rho)$, then the principal obtains an expected utility of, by Lemma 3.2:
$\frac{1}{T}\mathbb{E}\Big{[}\sum_{t=1}^{T}u(a^{t},\omega^{t})\Big{]}\,=\,U(\pi^{\varepsilon},\rho)\,\geq\min_{\rho\in\mathcal{R}_{\delta}(\pi^{\varepsilon})}U(\pi^{\varepsilon},\rho)\,\geq\,\underline{\mathrm{OBJ}}^{\mathcal{R}}\Big{(}\delta=\frac{\mathrm{CReg}(T)}{T}\Big{)}-\varepsilon$
in the learning model. Letting $\varepsilon\to 0$ proves the theorem. ∎
3.3 Agent’s No-Swap-Regret Learning: Upper Bound on Principal’s Utility
Theorem 3.3.
If the agent uses a contextual no-swap-regret learning algorithm, then the principal’s obtainable utility, even using time-varying strategies, is at most:
$\frac{1}{T}\mathbb{E}\Big{[}\sum_{t=1}^{T}u(x^{t},a^{t})\Big{]}\,\leq\,\overline{\mathrm{OBJ}}^{\mathcal{R}}\big{(}\frac{\mathrm{CSReg}(T)}{T}\big{)}.$
Proof.
Let $p_{s}^{t}=\Pr[s^{t}=s]=\mathbb{E}\big{[}\mathbbm{1}[s^{t}=s]\big{]}=\mathbb{E}[\pi_{s}^{t}]$ be the probability that signal $s\in S$ is sent in round $t$. Let $p_{a|s}^{t}=\Pr[a^{t}=a\,|\,s^{t}=s]$ be the probability that the agent takes action $a$ conditioning on signal $s^{t}=s$ is sent in round $t$. Let $d:S\times A\to A$ be any deviation function for the agent. The utility gain by deviation for the agent is upper bounded by the contextual swap-regret:
$\frac{\mathrm{CSReg}(T)}{T}\geq\frac{1}{T}\mathbb{E}\Big{[}\sum_{t=1}^{T}\Big{(}v(x^{t},d(s^{t},a^{t}))-v(x^{t},a^{t})\Big{)}\Big{]}$ (15)
$~{}=\frac{1}{T}\sum_{t=1}^{T}\sum_{s\in S}p_{s}^{t}\sum_{a\in A}p_{a|s}^{t}\mathbb{E}_{x^{t}_{s}|s^{t}=s}\Big{[}v(x^{t}_{s},d(s,a))-v(x^{t}_{s},a)\Big{]}$
$~{}=\frac{1}{T}\sum_{t=1}^{T}\sum_{s\in S}p_{s}^{t}\sum_{a\in A}p_{a|s}^{t}\Big{(}v(\mathbb{E}[x_{s}^{t}|s^{t}=s],d(s,a))-v(\mathbb{E}[x_{s}^{t}|s^{t}=s],a)\Big{)}\quad\quad\text{by linearity of $v(\cdot,a)$}$
$~{}=\sum_{s\in S}\sum_{a\in A}\frac{\sum_{j=1}^{T}p_{s}^{j}p_{a|s}^{j}}{T}\frac{1}{\sum_{j=1}^{T}p_{s}^{j}p_{a|s}^{j}}\sum_{t=1}^{T}p_{s}^{t}p_{a|s}^{t}\Big{(}v(\mathbb{E}[x_{s}^{t}|s^{t}=s],d(s,a))-v(\mathbb{E}[x_{s}^{t}|s^{t}=s],a)\Big{)}$
$~{}=\sum_{s\in S}\sum_{a\in A}\frac{\sum_{j=1}^{T}p_{s}^{j}p_{a|s}^{j}}{T}\bigg{[}v\Big{(}\tfrac{\sum_{t=1}^{T}p_{s}^{t}p_{a|s}^{t}\mathbb{E}[x_{s}^{t}|s^{t}=s]}{\sum_{j=1}^{T}p_{s}^{j}p_{a|s}^{j}},d(s,a)\Big{)}-v\Big{(}\tfrac{\sum_{t=1}^{T}p_{s}^{t}p_{a|s}^{t}\mathbb{E}[x_{s}^{t}|s^{t}=s]}{\sum_{j=1}^{T}p_{s}^{j}p_{a|s}^{j}},a\Big{)}\bigg{]}.$
Define $q_{s,a}=\frac{\sum_{j=1}^{T}p_{s}^{j}p_{a|s}^{j}}{T}$ and $y_{s,a}=\frac{\sum_{t=1}^{T}p_{s}^{t}p_{a|s}^{t}\mathbb{E}[x_{s}^{t}|s^{t}=s]}{\sum_{j=1}^{T}p_{s}^{j}p_{a|s}^{j}}\in\mathcal{X}$. Then the above is equal to
$=\sum_{s\in S}\sum_{a\in A}q_{s,a}\Big{[}v(y_{s,a},d(s,a))-v(y_{s,a},a)\Big{]}.$ (16)
We note that $\sum_{s\in S}\sum_{a\in A}q_{s,a}=\frac{\sum_{j=1}^{T}\sum_{s\in S}\sum_{a\in A}p_{s}^{j}p_{a|s}^{j}}{T}=1$, so $q$ is a probability distribution over $S\times A$. And note that
$\sum_{s,a\in S\times A}q_{s,a}y_{s,a}$ $~{}=~{}\sum_{s,a\in S\times A}\frac{1}{T}\sum_{t=1}^{T}p_{s}^{t}p_{a|s}^{t}\mathbb{E}[x_{s}^{t}|s^{t}=s]~{}=~{}\frac{1}{T}\sum_{t=1}^{T}\sum_{s\in S}p_{s}^{t}\mathbb{E}[x_{s}^{t}|s^{t}=s]$
$~{}=~{}\frac{1}{T}\sum_{t=1}^{T}\sum_{s\in S}\mathbb{E}\big{[}\mathbbm{1}[s^{t}=s]x_{s}^{t}\big{]}~{}=~{}\frac{1}{T}\sum_{t=1}^{T}\mathbb{E}\big{[}\sum_{s\in S}\mathbbm{1}[s^{t}=s]x_{s}^{t}\big{]}~{}=~{}\frac{1}{T}\sum_{t=1}^{T}\mathbb{E}\big{[}x^{t}\big{]}$
$~{}=~{}\frac{1}{T}\sum_{t=1}^{T}\mathbb{E}\big{[}\sum_{s\in S}\pi^{t}_{s}x_{s}^{t}\big{]}\in\mathcal{C}\quad\quad\text{because $\sum_{s\in S}\pi_{s}^{t}x_{s}^{t}\in\mathcal{C}$.}$
This means that $\pi^{\prime}=\{(q_{s,a},y_{s,a})\}_{(s,a)\in S\times A}$ defines a valid principal strategy with the larger signal space $S\times A$. Then, we note that (16) is the difference between the agent’s expected utility under principal strategy $\pi^{\prime}$ when responding using strategy $d:S\times A\to A$ and using the strategy that maps signal $(s,a)$ to action $a$. And (16) is upper bounded by $\frac{\mathrm{CSReg}(T)}{T}$ by (15):
(16)$=V(\pi^{\prime},d)-V(\pi^{\prime},(s,a)\mapsto a)\,\leq\,\tfrac{\mathrm{CSReg}(T)}{T},~{}~{}~{}\forall d:S\times A\to A.$ (17)
In particular, this holds when $d$ is the agent’s best-responding strategy. This means that the agent strategy $(s,a)\mapsto a$ is a $(\frac{\mathrm{CSReg}(T)}{T})$-best-response to $\pi^{\prime}$. So, the principal’s expected utility is upper bounded by the utility in the approximate-best-response model:
$\frac{1}{T}\mathbb{E}\Big{[}\sum_{t=1}^{T}u(x^{t},a^{t})\Big{]}$ $=\frac{1}{T}\sum_{t=1}^{T}\sum_{s\in S}p_{s}^{t}\sum_{a\in A}p_{a|s}^{t}v(\mathbb{E}[x_{s}^{t}|s^{t}=s],a)$
$=\sum_{s\in S}\sum_{a\in A}q_{s,a}u(y_{s,a},a)~{}=~{}U(\pi^{\prime},(s,a)\to a)~{}\leq~{}\overline{\mathrm{OBJ}}^{\mathcal{R}}\big{(}\tfrac{\mathrm{CSReg}(T)}{T}\big{)}.\qed$
3.4 Agent’s Mean-Based Learning: Exploitable by the Principal
Many no-regret (but not no-swap-regret) learning algorithms (e.g., MWU, FTPL, EXP-3) satisfy the following contextual mean-based property:
Definition 3.1 (Braverman et al. (2018)).
Let $\sigma_{s}^{t}(a)=\sum_{j\in[t]:s^{j}=s}v(\omega^{j},a)$ be the sum of historical utility of the receiver in the first $t$ rounds if he takes action $a$ when the signal/context is $s$. An algorithm is called $\gamma$-mean-based if: whenever $\exists a^{\prime}$ such that $\sigma_{s}^{t-1}(a)<\sigma_{s}^{t-1}(a^{\prime})-\gamma T$, the probability that the algorithm chooses action $a$ at round $t$ if the context is $s$ is $\Pr[a^{t}=a|s^{t}=s]<\gamma$, with $\gamma=o(1)$.
Theorem 3.4.
There exists a Bayesian persuasion instance where, as long as the receiver does $\gamma$-mean-based learning, the sender can obtain a utility significantly larger than $\overline{\mathrm{OBJ}}^{\mathcal{R}}(\gamma)$ and $U^{*}$.
Proof
The instance has $2$ states (A, B), $3$ actions (L, M, R), uniform prior $\mu_{0}(\rm{A})=\mu_{0}(\rm B)=0.5$, with the following utility matrices (left for sender’s, right for receiver’s):
$u(\omega,a)$ L M R $v(\omega,a)$ L M R
A 00 $-2$ $-2$ A $\sqrt{\gamma}$ $-1$ 00
B 00 00 $2$ B $-1$ $1$ 00
Claim 3.5.
In this instance, the optimal sender utility $U^{*}$ in the classic BP model is 00, and the approximate-best-response objective $\overline{\mathrm{OBJ}}^{\mathcal{R}}(\gamma)=O(\gamma)$.
Proof.
Recall that any signaling scheme decomposes the prior $\mu_{0}$ into multiple posteriors $\{\mu_{s}\}_{s\in S}$. If a posterior $\mu_{s}$ puts probability $>0.5$ to state B, then the receiver will take action M, which gives the sender a utility $\leq 0$; if the posterior $\mu_{s}$ puts probability $\leq 0.5$ to state B, then no matter what action the receiver takes, the sender’s expected utility on $\mu_{s}$ cannot be greater than 00. So, the sender’s expected utility is $\leq 0$ under any signaling scheme. An optimal signaling scheme is to reveal no information (keep $\mu_{s}=\mu_{0}$); the receiver takes R and the sender gets utility 00.
This instance satisfies the assumptions of Theorem 4.3, so $\overline{\mathrm{OBJ}}^{\mathcal{R}}(\gamma)\leq U^{*}+O(\gamma)=O(\gamma)$. ∎
Claim 3.6.
By doing the following, the sender can obtain utility $\approx\frac{1}{2}-O(\sqrt{\gamma})$ if the receiver is $\gamma$-mean-based learning:
  • in the first $T/2$ rounds: if the state is A, send signal 1; if the state is B, send 2.
  • in the remaining $T/2$ rounds, switch the scheme: if the state is A, send 2; if state is B, send 1.
Proof Sketch.
In the first $T/2$ rounds, the receiver finds that signal 1 corresponds to state A so he will take action L with high probability when signal 1 is sent; signal 2 corresponds to B so he will take action M with high probability. In this phase, the sender obtains utility $\approx 0$ per round. At the end of this phase, for signal 1, the receiver accumulates utility $\approx\frac{T}{2}\frac{1}{2}\sqrt{\gamma}=\frac{T}{4}\sqrt{\gamma}$ for action L. For signal 2, the receiver accumulates utility $\approx\frac{T}{2}\frac{1}{2}\cdot 1=\frac{T}{4}$ for action M.
In the remaining $T/2$ rounds, the following will happen:
  • For signal 1, the receiver finds that the state is now B, so the utility of action L decreases by 1 every time signal 1 is sent. Because the utility of L accumulated in the first phase was $\approx\frac{T}{4}\sqrt{\gamma}$, after $\approx\frac{T}{4}\sqrt{\gamma}$ rounds in second phase the utility of L should decrease to below 00, and the receiver will no longer play L (with high probability) at signal 1. The receiver will not play M at signal 1 in most of the second phase either, because there are more A states than B states at signal 1 historically. So, the receiver will play action R most times, roughly $\frac{T}{4}-\frac{T}{4}\sqrt{\gamma}$ rounds. This gives the sender a total utility of $\approx(\frac{T}{4}-\frac{T}{4}\sqrt{\gamma})\cdot 2=\frac{T}{2}-O(T\sqrt{\gamma})$.
  • For signal 2, the state is now A. But the receiver will continue to play action M in most times. This because: R has utility 0; L accumulated $\approx-\frac{T}{4}$ utility in the first phase, and only increases by $\sqrt{\gamma}$ per round in the second phase, so its accumulated utility is always negative; instead, M has accumulated $\frac{T}{4}$ utility in the first phase, and decreases by 1 every time signal $2$ is sent in the second phase, so its utility is positive until near the end. So, the receiver will play M. This gives the sender utility 00.
Summing up, the sender obtains total utility $\approx\frac{T}{2}-O(T\sqrt{\gamma})$ in these two phases, which is $\frac{1}{2}-O(\sqrt{\gamma})>0$ per round in average. ∎
4 Generalized Principal-Agent Problems with Approximate Best Response
Having presented the reduction from learning to approximate best response in Section 3, we now turn to study generalized principal-agent problems with approximate best response. We derive lower bounds on the maximin objectives $\underline{\mathrm{OBJ}}^{\mathcal{D}}(\delta)$, $\underline{\mathrm{OBJ}}^{\mathcal{R}}(\delta)$, and upper bounds on the maximax objectives $\overline{\mathrm{OBJ}}^{\mathcal{D}}(\delta)$, $\overline{\mathrm{OBJ}}^{\mathcal{R}}(\delta)$. We will show that, when the degree $\delta$ of agent’s approximate best response is small, all the four objectives will be close to the optimal principal objective $U^{*}$ in the best-response model, under some natural assumptions below.
Assumptions and notations
We make some innocuous assumptions. First, the agent has no weakly dominated action:
Assumption 4.1 (No Dominated Action).
An action $a_{0}\in A$ of the agent is weakly dominated if there exists a mixed action $\alpha^{\prime}\in\Delta(A\setminus\{a_{0}\})$ such that $v(x,\alpha^{\prime})=\mathbb{E}_{a\sim\alpha^{\prime}}[v(x,a)]\geq v(x,a_{0})$ for all $x\in\mathcal{X}$. We assume that the agent has no weakly dominated action.
Claim 4.1.
Assumption 4.1 implies: there exists a constant $G>0$ such that, for any agent action $a\in A$, there exists a principal decision $x\in\mathcal{X}$ such that $v(x,a)-v(x,a^{\prime})\geq G$ for every $a^{\prime}\in A\setminus\{a\}$.
The proof of this claim is in Appendix C.1.
The constant $G>0$ in Claim 4.1 is analogous to the concept of “inducibility gap” in Stackelberg games (Von Stengel and Zamir, 2004; Gan et al., 2023). In fact, Gan et al. (2023) show that, if the inducibility gap $G>\delta$, then the maximin approximate-best-response objective satisfies $\underline{\mathrm{OBJ}}^{\mathcal{D}}(\delta)\geq U^{*}-\frac{\delta}{G}$ in Stackelberg games. Our results will significantly generalize theirs to any generalized principal-agent problem, to randomized agent strategies, and to the maximax objectives $\overline{\mathrm{OBJ}}^{\mathcal{D}}(\delta)$, $\overline{\mathrm{OBJ}}^{\mathcal{R}}(\delta)$.
To present our results, we need to introduce a few more notions and assumptions. Let
$\mathrm{diam}(\mathcal{X};\|\cdot\|)=\max_{x_{1},x_{2}\in\mathcal{X}}\|x_{1}-x_{2}\|$ (18)
[…]
$\mathrm{dist}(\mathcal{C},\partial\mathcal{X})=\min_{c\in\mathcal{C},x\in\partial X}\|c-x\|$ (19)
be the distance from $\mathcal{C}$ to the boundary of $\mathcal{X}$. We assume that $\mathcal{C}$ is away from the boundary of $\mathcal{X}$:
Assumption 4.2 ($\mathcal{C}$ is in the interior of $\mathcal{X}$).
$\mathrm{dist}(\mathcal{C},\partial\mathcal{X})>0$.
Assumption 4.3 (Bounded and Lipschitz utility).
The principal’s utility function is bounded: $|u(x,a)|\leq B$, and $L$-Lipschitz in $x\in\mathcal{X}$: $|u(x_{1},a)-u(x_{2},a)|\leq L\|x_{1}-x_{2}\|$.
Main results
We now present the main results of this section: lower bounds on $\underline{\mathrm{OBJ}}^{X}(\delta)$ and upper bounds on $\overline{\mathrm{OBJ}}^{X}(\delta)$ in generalized principal-agent problems without and with constraints.
Theorem 4.2 (Without constraint).
For an unconstrained generalized principal-agent problem, under Assumptions 4.1 and 4.3, for $0\leq\delta<G$, we have
  • $\underline{\mathrm{OBJ}}^{\mathcal{D}}(\delta)\,\geq\,U^{*}-\mathrm{diam}(\mathcal{X})L\tfrac{\delta}{G}$.
  • $\underline{\mathrm{OBJ}}^{\mathcal{R}}(\delta)\,\geq\,U^{*}-2\sqrt{\tfrac{2BL}{G}\mathrm{diam}(\mathcal{X})\delta}~{}$ for $\delta<\frac{\mathrm{diam}(\mathcal{X})GL}{2B}$.
  • $\overline{\mathrm{OBJ}}^{\mathcal{D}}(\delta)\,\leq\,\overline{\mathrm{OBJ}}^{\mathcal{R}}(\delta)\,\leq\,U^{*}+\mathrm{diam}(\mathcal{X})L\tfrac{\delta}{G}$.
Theorem 4.3 (With constraint).
For a generalized principal-agent problem with the constraint $\sum_{s\in S}\pi_{s}x_{s}\in\mathcal{C}$, under Assumptions 4.1, 4.2 and 4.3, for $0\leq\delta<\frac{G\mathrm{dist}(\mathcal{C},\partial\mathcal{X})}{\mathrm{diam}(\mathcal{X})}$, we have
  • $\underline{\mathrm{OBJ}}^{\mathcal{D}}(\delta)\,\geq\,U^{*}-\big{(}\mathrm{diam}(\mathcal{X})L+2B\tfrac{\mathrm{diam}(\mathcal{X})}{\mathrm{dist}(\mathcal{C},\partial\mathcal{X})}\big{)}\tfrac{\delta}{G}$.
  • $\underline{\mathrm{OBJ}}^{\mathcal{R}}(\delta)\,\geq\,U^{*}-2\sqrt{\tfrac{2B}{G}\big{(}\mathrm{diam}(\mathcal{X})L+2B\tfrac{\mathrm{diam}(\mathcal{X})}{\mathrm{dist}(\mathcal{C},\partial\mathcal{X})}\big{)}\delta}$.
  • $\overline{\mathrm{OBJ}}^{\mathcal{D}}(\delta)\,\leq\,\overline{\mathrm{OBJ}}^{\mathcal{R}}(\delta)\,\leq\,U^{*}+\big{(}\mathrm{diam}(\mathcal{X})L+2B\tfrac{\mathrm{diam}(\mathcal{X})}{\mathrm{dist}(\mathcal{C},\partial\mathcal{X})}\big{)}\tfrac{\delta}{G}$.
The expression “$\tfrac{\mathrm{diam}(\mathcal{X})}{\mathrm{dist}(\mathcal{C},\partial\mathcal{X})}\delta$” suggests that $\frac{1}{\mathrm{dist}(\mathcal{C},\partial\mathcal{X})}$ is similar to a “condition number” (Renegar, 1994) that quantifies the “stability” of the principal-agent problem against the agent’s approximate-best-responding behavior. When $\mathrm{dist}(\mathcal{C},\partial\mathcal{X})$ is larger ($\mathcal{C}$ is further away from the boundary of $\mathcal{X}$), the condition number is smaller, the problem is more stable, and the $\delta$-best-response objectives $\underline{\mathrm{OBJ}}^{X}(\delta)$, and $\overline{\mathrm{OBJ}}^{X}(\delta)$ are closer to the best-response objective $U^{*}$.
Lower bounds for $\underline{\mathrm{OBJ}}^{\mathcal{D}}(\delta)$ imply lower bounds for $\underline{\mathrm{OBJ}}^{\mathcal{R}}(\delta)$
The lower bounds for $\underline{\mathrm{OBJ}}^{\mathcal{D}}(\delta)$ in Theorem 4.2 and 4.3 actually imply the lower bounds for $\underline{\mathrm{OBJ}}^{\mathcal{R}}(\delta)$. This is one step in the proofs of the two theorems, so we present it here. We use the following lemma (proved in Appendix C.2):
Lemma 4.4.
For any $\delta\geq 0,\Delta>0$, $\underline{\mathrm{OBJ}}^{\mathcal{R}}(\delta)\geq\underline{\mathrm{OBJ}}^{\mathcal{D}}(\Delta)-\tfrac{2B\delta}{\Delta}$.
Using Lemma 4.4 with $\Delta=\sqrt{\frac{2BG\delta}{\mathrm{diam}(\mathcal{X})L}}$ and the bound for $\underline{\mathrm{OBJ}}^{\mathcal{D}}(\Delta)$ in Theorem 4.2, we obtain:
$\underline{\mathrm{OBJ}}^{\mathcal{R}}(\delta)\geq\underline{\mathrm{OBJ}}^{\mathcal{D}}(\Delta)-\tfrac{2B\delta}{\Delta}$ $\geq U^{*}-\mathrm{diam}(\mathcal{X})L\tfrac{\Delta}{G}-\tfrac{2B\delta}{\Delta}=U^{*}-2\sqrt{\tfrac{2BL}{G}\mathrm{diam}(\mathcal{X})\delta},$
which gives the lower bound for $\underline{\mathrm{OBJ}}^{\mathcal{R}}(\delta)$ in Theorem 4.2.
Using Lemma 4.4 with $\Delta=\sqrt{\frac{2BG\delta}{L\mathrm{diam}(\mathcal{X})+2B\tfrac{\mathrm{diam}(\mathcal{X})}{\mathrm{dist}(\mathcal{C},\partial\mathcal{X})}}}$ and the bound for $\underline{\mathrm{OBJ}}^{\mathcal{D}}(\Delta)$ in Theorem 4.3,
$\underline{\mathrm{OBJ}}^{\mathcal{R}}(\delta)\geq\underline{\mathrm{OBJ}}^{\mathcal{D}}(\Delta)-\tfrac{2B\delta}{\Delta}$ $\geq U^{*}-\big{(}\mathrm{diam}(\mathcal{X})L+2B\tfrac{\mathrm{diam}(\mathcal{X})}{\mathrm{dist}(\mathcal{C},\partial\mathcal{X})}\big{)}\tfrac{\Delta}{G}-\tfrac{2B\delta}{\Delta}$
$=U^{*}-2\sqrt{\tfrac{2B}{G}\big{(}\mathrm{diam}(\mathcal{X})L+2B\tfrac{\mathrm{diam}(\mathcal{X})}{\mathrm{dist}(\mathcal{C},\partial\mathcal{X})}\big{)}\delta}.$
This proves the lower bound for $\underline{\mathrm{OBJ}}^{\mathcal{R}}(\delta)$ in Theorem 4.3.
The bound $\underline{\mathrm{OBJ}}^{\mathcal{R}}(\delta)\geq U^{*}-O(\sqrt{\delta})$ is tight.
We note that, in Theorem 4.2 and 4.3, the maximin objective with randomized agent strategies is bounded by $\underline{\mathrm{OBJ}}^{\mathcal{R}}(\delta)\geq U^{*}-O(\sqrt{\delta})$ while the objective with deterministic agent strategies is bounded by $\underline{\mathrm{OBJ}}^{\mathcal{D}}(\delta)\geq U^{*}-O(\delta)$. This is not because our analysis is not tight. In fact, the tight bound for $\underline{\mathrm{OBJ}}^{\mathcal{R}}(\delta)$ is $U^{*}-\Theta(\sqrt{\delta})$, with a squared root dependency on $\delta$. We prove this by giving an example where $\underline{\mathrm{OBJ}}^{\mathcal{R}}(\delta)\leq U^{*}-\Omega(\sqrt{\delta})$. Consider the following classical Bayesian persuasion example with two states and two actions:
Example 4.1.
There are two states $\Omega=\{\mathrm{Good},\mathrm{Bad}\}$, two actions $A=\{a,b\}$. The sender wants the receiver to take action $a$ regardless of the state. Action $a$ gives the receiver utility $1$ at $\mathrm{Good}$ state and $-1$ at $\mathrm{Bad}$ state; action $b$ gives utility 00. The prior is represented by a real number $\mu_{0}\in[0,1]$ denoting the probability of $\mathrm{Good}$ state. Assume $\mu_{0}<\frac{1}{2}$, so the receiver will take action $b$ by default.
sender $a$ $b$ receiver $a$ $b$
$\mathrm{Good}$ $1$ 00 $\mathrm{Good}$ $1$ 00
$\mathrm{Bad}$ $1$ 00 $\mathrm{Bad}$ $-1$ 00
In this example, for $\delta<\frac{\mu_{0}}{2}$, $\underline{\mathrm{OBJ}}^{\mathcal{R}}(\delta)\leq U^{*}-2\sqrt{2\mu_{0}\delta}+\delta=U^{*}-\Omega(\sqrt{\delta})$. (The proof is in Appendix C.3.)
4.1 Proof of Theorem 4.2 and 4.3
Since we have shown that the lower bounds for $\underline{\mathrm{OBJ}}^{\mathcal{D}}(\delta)$ imply the lower bounds for $\underline{\mathrm{OBJ}}^{\mathcal{R}}(\delta)$, we only need to prove the lower bounds for $\underline{\mathrm{OBJ}}^{\mathcal{D}}(\delta)$ and the upper bounds for $\overline{\mathrm{OBJ}}^{\mathcal{R}}(\delta)$.
Lemma 4.5.
In an unconstrained generalized principal-agent problem, $\underline{\mathrm{OBJ}}^{\mathcal{D}}(\delta)\geq U^{*}-\mathrm{diam}(\mathcal{X})L\tfrac{\delta}{G}$.
With the constraint $\sum_{s\in S}\pi_{s}x_{s}\in\mathcal{C}$, $\underline{\mathrm{OBJ}}^{\mathcal{D}}(\delta)\geq U^{*}-\big{(}\mathrm{diam}(\mathcal{X})L+2B\tfrac{\mathrm{diam}(\mathcal{X})}{\mathrm{dist}(\mathcal{C},\partial\mathcal{X})}\big{)}\tfrac{\delta}{G}$.
Lemma 4.6.
In an unconstrained generalized principal-agent problem, $\overline{\mathrm{OBJ}}^{\mathcal{R}}(\delta)\leq U^{*}+\mathrm{diam}(\mathcal{X})L\tfrac{\delta}{G}$.
With the constraint $\sum_{s\in S}\pi_{s}x_{s}\in\mathcal{C}$, $\overline{\mathrm{OBJ}}^{\mathcal{R}}(\delta)\leq U^{*}+\big{(}\mathrm{diam}(\mathcal{X})L+2B\tfrac{\mathrm{diam}(\mathcal{X})}{\mathrm{dist}(\mathcal{C},\partial\mathcal{X})}\big{)}\tfrac{\delta}{G}$.
Lemma 4.5 and 4.6 immediately prove Theorem 4.2 and 4.3.
We prove Lemma 4.6 here and Lemma 4.5 in Appendix C.4. The main idea is the following: let $(\pi,\rho)$ be a pair of principal’s strategy and agent’s $\delta$-best-responding strategy that gives the principal a high utility. We perturb the principal’s strategy $\pi$ slightly to be a strategy $\pi^{\prime}$ for which $\rho$ is best-responding (such a perturbation is possible due to Assumption 4.1). Since $\rho$ is best-responding to $\pi^{\prime}$, the pair $(\pi^{\prime},\rho)$ cannot give the principal a higher utility than $U^{*}$ (which is the optimal principal utility under the best-response model). This means that the original pair $(\pi,\rho)$ cannot give the principal a utility much higher than $U^{*}$.
Proof of Lemma 4.6.
Let $\pi$ be a principal strategy and $\rho\in\mathcal{R}_{\delta}(\pi)$ be a $\delta$-best-responding randomized strategy of the agent. The principal strategy $\pi$ consists of pairs $\{(\pi_{s},x_{s})\}_{s\in S}$ with
$\sum_{s\in S}\pi_{s}x_{s}=:\mu_{0}\in\mathcal{C}.$ (20)
At signal $s$, the agent takes action $a$ with probability $\rho(a|s)$. Let $\delta_{s,a}$ be the “suboptimality” of action $a$ with respect to $x_{s}$:
$\delta_{s,a}=\max_{a^{\prime}\in A}\big{\{}v(x_{s},a^{\prime})-v(x_{s},a)\big{\}}.$ (21)
By Claim 4.1, for action $a$ there exists $y_{a}\in\mathcal{X}$ such that $v(y_{a},a)-v(y_{a},a^{\prime})\geq G$ for any $a^{\prime}\neq a$. Let $\theta_{s,a}=\frac{\delta_{s,a}}{G+\delta_{s,a}}\in[0,1]$ and let $\tilde{x}_{s,a}$ be the convex combination of $x_{s}$ and $y_{a}$ with weights $1-\theta_{s,a},\theta_{s,a}$:
$\tilde{x}_{s,a}=(1-\theta_{s,a})x_{s}+\theta_{s,a}y_{a}.$ (22)
Claim 4.7.
We have two useful claims regarding $\tilde{x}_{s,a}$ and $\theta_{s,a}$:
  • (1)
    $a$ is an optimal action for the agent with respect to $\tilde{x}_{s,a}$: $v(\tilde{x}_{s,a},a)-v(\tilde{x}_{s,a},a^{\prime})\geq 0,\forall a^{\prime}\in A$.
  • (2)
    $\sum_{s\in S}\sum_{a\in A}\pi_{s}\rho(a|s)\theta_{s,a}\leq\frac{\delta}{G}$.
Proof.
(1) For any $a^{\prime}\neq a$, by the definition of $\tilde{x}_{s,a}$ and $\theta_{s,a}$,
$v(\tilde{x}_{s,a},a)-v(\tilde{x}_{s,a},a^{\prime})$ $~{}=~{}(1-\theta_{s,a})\big{[}v(x_{s},a)-v(x_{s},a^{\prime})\big{]}~{}+~{}\theta_{s,a}\big{[}v(y_{a},a)-v(y_{a},a^{\prime})\big{]}$
$~{}\geq~{}(1-\theta_{s,a})(-\delta_{s,a})~{}+~{}\theta_{s,a}G~{}=~{}\tfrac{G}{G+\delta_{s,a}}(-\delta_{s,a})+\tfrac{\delta_{s,a}}{G+\delta_{s,a}}G~{}=~{}0.$
(2) By the condition that $\rho$ is a $\delta$-best-response to $\pi$, we have
$\delta~{}$ $\geq~{}\max_{\rho^{*}:S\to A}V(\pi,\rho^{*})-V(\pi,\rho)~{}=~{}\sum_{s\in S}\pi_{s}\Big{(}\max_{a^{\prime}\in A}\big{\{}v(x_{s},a^{\prime})\big{\}}-\sum_{a\in A}\rho(a|s)v(x_{s},a)\Big{)}$
$=~{}\sum_{s\in S}\sum_{a\in A}\pi_{s}\rho(a|s)\max_{a^{\prime}\in A}\big{\{}v(x_{s},a^{\prime})-v(x_{s},a)\big{\}}~{}=~{}\sum_{s\in S}\sum_{a\in A}\pi_{s}\rho(a|s)\delta_{s,a}.$
So, $\sum_{s\in S}\sum_{a\in A}\pi_{s}\rho(a|s)\theta_{s,a}=\sum_{s\in S}\sum_{a\in A}\pi_{s}\rho(a|s)\frac{\delta_{s,a}}{G+\delta_{s,a}}\leq\sum_{s\in S}\sum_{a\in A}\pi_{s}\rho(a|s)\frac{\delta_{s,a}}{G}\leq\frac{\delta}{G}$. ∎
We let $\mu^{\prime}$ be the convex combination of $\{\tilde{x}_{s,a}\}_{s,a\in S\times A}$ with weights $\{\pi_{s}\rho(a|s)\}_{s,a\in S\times A}$:
$\mu^{\prime}=\sum_{s,a\in S\times A}\pi_{s}\rho(a|s)\tilde{x}_{s,a}.$ (23)
Note that $\mu^{\prime}$ might not satisfy the constraint $\mu^{\prime}\in\mathcal{C}$. So, we want to find another vector $z\in\mathcal{X}$ and a coefficient $\eta\in[0,1]$ such that
$(1-\eta)\mu^{\prime}+\eta z\in\mathcal{C}.$ (24)
(If $\mu^{\prime}$ already satisfies $\mu^{\prime}\in\mathcal{C}$, then let $\eta=0$.) To do this, we consider the ray pointing from $\mu^{\prime}$ to $\mu_{0}$: $\{\mu^{\prime}+t(\mu_{0}-\mu^{\prime})\mid t\geq 0\}$. Let $z$ be the intersection of the ray with the boundary of $\mathcal{X}$:
$z=\mu^{\prime}+t^{*}(\mu_{0}-\mu^{\prime}),\quad\quad t^{*}=\operatorname*{arg\,max}\{t\geq 0\mid\mu^{\prime}+t(\mu_{0}-\mu^{\prime})\in\mathcal{X}\}.$
Then, rearranging $z=\mu^{\prime}+t^{*}(\mu_{0}-\mu^{\prime})$, we get
$\tfrac{1}{t^{*}}(z-\mu^{\prime})=\mu_{0}-\mu^{\prime}\quad\;\Longleftrightarrow\;\quad(1-\tfrac{1}{t^{*}})\mu^{\prime}+\tfrac{1}{t^{*}}z=\mu_{0}\in\mathcal{C},$
which satisfies (24) with $\eta=\frac{1}{t^{*}}$. We then give an upper bound on $\eta=\frac{1}{t*}$:
Claim 4.8.
$\eta=\tfrac{1}{t^{*}}\leq\tfrac{\mathrm{diam}(\mathcal{X})}{\mathrm{dist}(\mathcal{C},\partial\mathcal{X})}\tfrac{\delta}{G}$. (See proof in Appendix C.5.)
The convex combinations (24) (23) define a new principal strategy $\pi^{\prime}$ (with $|S|\times|A|+1$ signals) consisting of $\tilde{x}_{s,a}$ with probability $(1-\eta)\pi_{s}\rho(a|s)$ and $z$ with probability $\eta$. Consider the following deterministic agent strategy $\rho^{\prime}$ in response to $\pi^{\prime}$: for $\tilde{x}_{s,a}$, take action $\rho^{\prime}(\tilde{x}_{s,a})=a$; for $z$, take any action that is optimal for $z$. We note that $\rho^{\prime}$ is a best-response to $\pi^{\prime}$, $\rho^{\prime}\in\mathcal{R}_{0}(\pi^{\prime})$, because, according to Claim 4.7, $a$ is an optimal action with respect to $\tilde{x}_{s,a}$.
Then, consider the principal’s utility under $\pi^{\prime}$ and $\rho^{\prime}$:
$U(\pi^{\prime},\rho^{\prime})$ =(24),(23)${{=}}(1-\eta)\sum_{s\in S}\sum_{a\in A}\pi_{s}\rho(a|s)u(\tilde{x}_{s,a},\rho^{\prime}(\tilde{x}_{s,a}))~{}+~{}\eta u(z,\rho^{\prime}(z))$ ) )
$\geq~{}(1-\eta)\sum_{s\in S}\sum_{a\in A}\pi_{s}\rho(a|s)u(\tilde{x}_{s,a},a)~{}-~{}\eta B$
$\geq~{}(1-\eta)\sum_{s\in S}\sum_{a\in A}\pi_{s}\rho(a|s)\Big{(}u(x_{s},a)-L\underbrace{\|\tilde{x}_{s}-x_{s}\|}_{=\|\theta_{s,a}(y_{a}-x_{s})\|\leq\theta_{s,a}\mathrm{diam}(\mathcal{X})}\Big{)}~{}-~{}\eta B$
$\geq~{}(1-\eta)U(\pi,\rho)~{}-~{}L\mathrm{diam}(\mathcal{X})\sum_{s\in S}\sum_{a\in A}\pi_{s}\rho(a|s)\theta_{s,a}~{}-~{}\eta B$
(Claim 4.7) $\geq~{}U(\pi,\rho)-L\mathrm{diam}(\mathcal{X})\tfrac{\delta}{G}-2\eta B$
(Claim 4.8) $\geq~{}U(\pi,\rho)-\big{(}L\mathrm{diam}(\mathcal{X})+2B\tfrac{\mathrm{diam}(\mathcal{X})}{\mathrm{dist}(\mathcal{C},\partial\mathcal{X})}\big{)}\tfrac{\delta}{G}.$
Rearranging, $U(\pi,\rho)\leq U(\pi^{\prime},\rho^{\prime})+\big{(}L\mathrm{diam}(\mathcal{X})+2B\tfrac{\mathrm{diam}(\mathcal{X})}{\mathrm{dist}(\mathcal{C},\partial\mathcal{X})}\big{)}\tfrac{\delta}{G}$. Note that this argument holds for any pair $(\pi,\rho)$ that satisfies $\rho\in\mathcal{R}_{\delta}(\pi)$. And recall that $\rho^{\prime}\in\mathcal{R}_{0}(\pi^{\prime})$. So, we conclude that
$\overline{\mathrm{OBJ}}^{\mathcal{R}}(\delta)~{}=~{}\max_{\pi}\max_{\rho\in\mathcal{R}_{\delta}(\pi)}U(\pi,\rho)$ $~{}\leq~{}\max_{\pi^{\prime}}\max_{\rho^{\prime}\in\mathcal{R}_{0}(\pi)}U(\pi^{\prime},\rho^{\prime})+\big{(}L\mathrm{diam}(\mathcal{X};\ell_{1})+2B\tfrac{\mathrm{diam}(\mathcal{X})}{\mathrm{dist}(\mathcal{C},\partial\mathcal{X})}\big{)}\tfrac{\delta}{G}$
$~{}=~{}U^{*}+\big{(}L\mathrm{diam}(\mathcal{X};\ell_{1})+2B\tfrac{\mathrm{diam}(\mathcal{X})}{\mathrm{dist}(\mathcal{C},\partial\mathcal{X})}\big{)}\tfrac{\delta}{G}.$
This proves the case with the constraint $\sum_{s\in S}\pi_{s}x_{s}\in\mathcal{C}$.
The case without $\sum_{s\in S}\pi_{s}x_{s}\in\mathcal{C}$ is proved by letting $\eta=0$ in the above argument. ∎
5 Applications to Specific Principal-Agent Problems
In this section, we apply the general results in Section 3 and 4 to specific principal-agent problems to derive some concrete results for Bayesian persuasion, Stackelberg games, and contract design with learning agents. Our result for Bayesian persuasion is new, while our results for Stackelberg games and contract design refine Deng et al. (2019) and Guruganesh et al. (2024).
5.1 Bayesian Persuasion
As noted in Section 2, Bayesian persuasion is a generalized principal-agent problem with constraint
$\sum_{s\in S}\pi_{s}x_{s}\in\mathcal{C}=\{\mu_{0}\}$
where each $x_{s}=\mu_{s}=(\mu_{s}(\omega))_{\omega\in\Omega}\in\mathcal{X}=\Delta(\Omega)$ is a posterior belief. Suppose the principal’s utility is bounded: $|u(\omega,a)|\leq B$. Then, the principal’s utility function $u(\mu_{s},a)=\sum_{\omega\in\Omega}\mu_{s}(\omega)u(\omega,a)$ is $(L=B)$-Lipschitz (under $\ell_{1}$-norm):
$|u(\mu_{1},a)-u(\mu_{2},a)|\leq\sum_{\omega\in\Omega}|\mu_{1}(\omega)-\mu_{2}(\omega)|\cdot|u(\omega,a)\big{|}\leq\|\mu_{1}-\mu_{2}\|_{1}B,$ (25)
so Assumption 4.3 is satisfied. Suppose the prior $\mu_{0}$ has positive probability for every $\omega\in\Omega$, and let
$p_{0}=\min_{\omega\in\Omega}\mu_{0}(\omega)>0.$
Then, we have the distance
$\mathrm{dist}(\mathcal{C},\partial X)=\min\big{\{}\|\mu_{0}-\mu\|_{1}:\mu\in\Delta(\Omega)\text{ s.t. }\mu(\omega)=0\text{ for some $\omega\in\Omega$}\big{\}}\geq p_{0}>0,$
so Assumption 4.2 is satisfied. The diameter satisfies, because $\|\mu\|_{1}=1$ for $\mu\in\Delta(\Omega)$,
$\mathrm{diam}(\mathcal{X};\ell_{1})=\max_{\mu_{1},\mu_{2}\in\Delta(\Omega)}\|\mu_{1}-\mu_{2}\|_{1}\leq 2.$
Finally, we assume Assumption 4.1 (no dominated action for the agent). Then, Theorem 4.3 gives bounds on the approximate-best-response objectives in Bayesian persuasion:
Corollary 5.1 (Bayesian persuasion with approximate best response).
For $0\leq\delta<\frac{Gp_{0}}{2}$,
  • $\underline{\mathrm{OBJ}}^{\mathcal{D}}(\delta)\,\geq\,U^{*}-2B(1+\tfrac{2}{p_{0}})\tfrac{\delta}{G}$.
  • $\underline{\mathrm{OBJ}}^{\mathcal{R}}(\delta)\,\geq\,U^{*}-4B\sqrt{(1+\tfrac{2}{p_{0}})\tfrac{\delta}{G}}$.
  • $\overline{\mathrm{OBJ}}^{\mathcal{D}}(\delta)\,\leq\,\overline{\mathrm{OBJ}}^{\mathcal{R}}(\delta)\,\leq\,U^{*}+2B(1+\tfrac{2}{p_{0}})\tfrac{\delta}{G}$.
Further applying Theorem 3.1 and 3.3, we obtain the central result of our main problem, persuasion with a learning agent:
Corollary 5.2 (Persuasion with a learning agent).
Suppose $T$ is sufficiently large such that $\frac{\mathrm{CReg}(T)}{T}<\frac{Gp_{0}}{2}$ and $\frac{\mathrm{CSReg}(T)}{T}<\frac{Gp_{0}}{2}$, then
  • with a contextual no-regret learning agent, the principal can obtain utility at least
    $\frac{1}{T}\mathbb{E}\big{[}\sum_{t=1}^{T}u(x^{t},a^{t})\big{]}\,\geq\,\underline{\mathrm{OBJ}}^{\mathcal{R}}\big{(}\tfrac{\mathrm{CReg}(T)}{T}\big{)}\,\geq\,U^{*}-4B\sqrt{(1+\tfrac{2}{p_{0}})\tfrac{1}{G}}\sqrt{\tfrac{\mathrm{CReg}(T)}{T}}$ (26)
    using a fixed signaling schemes in all rounds.
  • with a contextual no-swap-regret learning agent, the principal’s obtainable utility is at most
    $\frac{1}{T}\mathbb{E}\big{[}\sum_{t=1}^{T}u(x^{t},a^{t})\big{]}\,\leq\,\overline{\mathrm{OBJ}}^{\mathcal{D}}\big{(}\tfrac{\mathrm{CSReg}(T)}{T}\big{)}\,\leq\,U^{*}+2B(1+\tfrac{2}{p_{0}})\tfrac{1}{G}\tfrac{\mathrm{CSReg}(T)}{T}$ (27)
    even using time-varying signaling schemes.
The result (27) is particularly interesting. Recall (from Section 2.3) that the learning agent actually chooses strategy before the principal, so the principal can best respond to the agent’s strategy. Nevertheless, (27) shows that the principal cannot do much better than the classic model where the principal moves first. This is a special property of no-swap-regret learning. If the agent does no-regret but not no-swap-regret learning, then the principal can do significantly better than $U^{*}$ in some instances, as we showed in Section 3.4.
5.2 Stackelberg Games
In a Stackelberg game, the principal (leader), having a finite action set $B$, first commits to a mixed strategy $x=(x_{(b)})_{b\in B}\in\Delta(B)$, which is a distribution over actions. So the principal’s decision space $\mathcal{X}$ is $\Delta(B)$. The agent (follower) then takes an action $a\in A$ in response to $x$. The (expected) utilities for the two players are $u(x,a)=\sum_{b\in B}x_{(b)}u(b,a)$ and $v(x,a)=\sum_{b\in B}x_{(b)}u(b,a)$. The signal $s$ can (but not necessarily) be an action that the principal recommends the agent to take.
Assume bounded utility $|u(b,a)|\leq B$. Then, the principal’s utility function is bounded in $[-B,B]$ and $(L=B)$-Lipschitz (by a similar argument as (25)). The diameter $\mathrm{diam}(\mathcal{X})=\max_{x_{1},x_{2}\in\Delta(B)}\|x_{1}-x_{2}\|_{1}\leq 2$. Applying the theorem for unconstrained generalized principal-agent problems (Theorem 4.2) and the theorems for learning agent (Theorem 3.1 and 3.3), we obtain:
Corollary 5.3 (Stackelberg game with a learning agent).
Suppose $T$ is sufficiently large such that $\frac{\mathrm{CReg}(T)}{T}<G$ and $\frac{\mathrm{CSReg}(T)}{T}<G$, then:
  • with a contextual no-regret learning agent, the principal can obtain utility $\frac{1}{T}\mathbb{E}\big{[}\sum_{t=1}^{T}u(x^{t},a^{t})\big{]}\geq\underline{\mathrm{OBJ}}^{\mathcal{R}}\big{(}\frac{\mathrm{CReg}(T)}{T}\big{)}\geq U^{*}-\frac{4B}{\sqrt{G}}\sqrt{\frac{\mathrm{CReg}(T)}{T}}$ using a fixed strategy in all rounds.
  • with a contextual no-swap-regret learning agent, the principal cannot obtain utility more than $\frac{1}{T}\mathbb{E}\big{[}\sum_{t=1}^{T}u(x^{t},a^{t})\big{]}\leq\overline{\mathrm{OBJ}}^{\mathcal{D}}\big{(}\frac{\mathrm{CSReg}(T)}{T}\big{)}\leq U^{*}+\tfrac{2B}{G}\frac{\mathrm{CSReg}(T)}{T}$ even using time-varying strategies.
The conclusion that the principal can obtain utility at least $U^{*}-o(1)$ against a no-regret learning agent and no more than $U^{*}+o(1)$ against a no-swap-regret agent in Stackelberg games was proved by Deng et al. (2019). Our Corollary 5.3 reproduces this conclusion and moreover provides bounds on the $o(1)$ terms, namely, $U^{*}-O(\sqrt{\frac{\mathrm{CReg}(T)}{T}})$ and $U^{*}+O(\frac{\mathrm{CSReg}(T)}{T})$. This demonstrates the generality and usefulness of our framework.
5.3 Contract Design
In contract design, there is a finite outcome space $O=\{r_{1},\ldots,r_{d}\}$ where each $r_{i}\in{\mathbb{R}}$ is a monetary reward to the principal. When the agent takes action $a\in A$, outcome $r_{i}$ will happen with probability $p_{ai}\geq 0$, $\sum_{i=1}^{d}p_{ai}=1$. The principal cannot observe the action taken by the agent but can observe the realized outcome. The principal’s decision space $\mathcal{X}$ is the set of contracts, where a contract $x=(x_{(i)})_{i=1}^{d}\in[0,+\infty]^{d}$ is a vector that specifies the payment to the agent for each possible outcome. So, if the agent takes action $a$ under contract $x$, the principal obtains expected utility
$u(x,a)=\sum_{i=1}^{d}p_{ai}(r_{i}-x_{(i)})$
and the agent obtains $v(x,a)=\sum_{i=1}^{d}p_{ai}x_{(i)}$. The signal $s$ can (but not necessarily) be an action that the principal recommends the agent to take. The principal’s decision space $\mathcal{X}\subseteq[0,+\infty]^{d}$ in contract design, however, may be unbounded and violate the requirement of bounded diameter $\mathrm{diam}(\mathcal{X})$ that we need. We have two remedies for this.
The first remedy is to require the principal’s payment to the agent be upper bounded by some constant $P<+\infty$, so $0\leq x_{(i)}\leq P$ and $\mathcal{X}=[0,P]^{d}$. Under this requirement and the assumption of bounded reward $|r_{i}|\leq R$, the principal’s utility becomes bounded by $|u(x,a)|\leq\sum_{i=1}^{d}p_{ai}(R+P)=R+P=B$ and $(L=1)$-Lipschitz under $\ell_{\infty}$-norm:
$|u(x_{1},a)-u(x_{2},a)|=\big{|}\sum_{i=1}^{d}p_{ai}(x_{1(i)}-x_{2(i)})\big{|}\leq\max_{i=1}^{d}|x_{1(i)}-x_{2(i)}|\sum_{i=1}^{d}p_{ai}=\|x_{1}-x_{2}\|_{\infty}.$ (28)
And the diameter of $\mathcal{X}$ is bounded by (under $\ell_{\infty}$-norm)
$\mathrm{diam}(\mathcal{X};\ell_{\infty})=\max_{x_{1},x_{2}\in\mathcal{X}}\|x_{1}-x_{2}\|_{\infty}=\max_{x_{1},x_{2}\in[0,P]^{d}}\max_{i=1}^{d}|x_{1(i)}-x_{2(i)}|\leq P.$ (29)
Now, we can apply the theorem for unconstrained generalized principal-agent problems (Theorem 4.2) and the theorems for learning agent (Theorem 3.1 and Theorem 3.3) to obtain:
Corollary 5.4 (Contract design (with bounded payment) with a learning agent).
Suppose $T$ is sufficiently large such that $\frac{\mathrm{CReg}(T)}{T}<\frac{PG}{2(R+P)}$ and $\frac{\mathrm{CSReg}(T)}{T}<G$, then:
  • with a contextual no-regret learning agent, the principal can obtain utility at least $\frac{1}{T}\mathbb{E}\big{[}\sum_{t=1}^{T}u(x^{t},a^{t})\big{]}$ $\geq\underline{\mathrm{OBJ}}^{\mathcal{R}}\big{(}\frac{\mathrm{CReg}(T)}{T}\big{)}\geq U^{*}-2\sqrt{\tfrac{2(R+P)P}{G}}\sqrt{\frac{\mathrm{CReg}(T)}{T}}$ using a fixed contract in all rounds.
  • with contextual a no-swap-regret learning agent, the principal cannot obtain utility more than $\frac{1}{T}\mathbb{E}\big{[}\sum_{t=1}^{T}u(x^{t},a^{t})\big{]}\leq\overline{\mathrm{OBJ}}^{\mathcal{D}}\big{(}\frac{\mathrm{CSReg}(T)}{T}\big{)}\leq U^{*}+\tfrac{P}{G}\frac{\mathrm{CSReg}(T)}{T}$ even using time-varying contracts.
The second remedy is to write contract design as a generalized principal-agent problem in another way. Let $\tilde{x}=(\tilde{x}_{(a)})_{a\in A}\in[0,+\infty]^{|A|}$ be a vector recording the expected payment from the principal to the agent for each action $a\in A$:
$\tilde{x}_{(a)}=\sum_{i=1}^{d}p_{ai}x_{(i)}.$ (30)
And let $\tilde{r}_{(a)}$ be the expected reward of action $a$, $\tilde{r}_{(a)}=\sum_{i=1}^{d}p_{ai}r_{i}$. Then, the principal and the agent’s utility can be rewritten as functions of $\tilde{x}$ and $a$:
$u(\tilde{x},a)=\tilde{r}_{(a)}-\tilde{x}_{(a)},\quad\quad v(\tilde{x},a)=\tilde{x}_{(a)},$ (31)
which are linear in $\tilde{x}\in\tilde{\mathcal{X}}$. Assuming bounded reward $|\tilde{r}_{(a)}|\leq R$, we can without loss of generality assume that the expected payment $\tilde{x}_{(a)}$ is bounded by $R$ as well, because otherwise the principal will get negative utility. So, the principal’s decision space can be restricted to
$\tilde{\mathcal{X}}=\big{\{}\tilde{x}\mid\text{$\exists$ $x\in[0,+\infty]^{d}$such that $\tilde{x}_{(a)}=\sum_{i=1}^{d}p_{ai}x_{(i)}$ for every $a\in A$}\big{\}}\,\cap\,[0,R]^{|A|},$ (32)
which is convex and has bounded diameter (under $\ell_{\infty}$ norm)
$\mathrm{diam}(\tilde{\mathcal{X}};\ell_{\infty})\leq\mathrm{diam}([0,R]^{|A|};\ell_{\infty})=R.$ (33)
The utility function $u(\tilde{x},a)$ is bounded by $2R$ and $(L=1)$-Lipschitz (under $\ell_{\infty}$ norm):
$|u(\tilde{x}_{1},a)-u(\tilde{x}_{2},a)|=|\tilde{x}_{1(a)}-\tilde{x}_{2(a)}|\leq\max_{a\in A}|\tilde{x}_{1(a)}-\tilde{x}_{2(a)}|=\|\tilde{x}_{1}-\tilde{x}_{2}\|_{\infty}.$ (34)
Thus, we can apply the theorem for unconstrained generalized principal-agent problems (Theorem 4.2) and the theorems for learning agent (Theorem 3.1 and Theorem 3.3) to obtain:
Corollary 5.5 (Contract design with a learning agent).
Suppose $T$ is sufficiently large such that $\frac{\mathrm{CReg}(T)}{T}<\frac{G}{2}$ and $\frac{\mathrm{CSReg}(T)}{T}<G$, then:
  • with a contextual no-regret learning agent, the principal can obtain utility at least $\frac{1}{T}\mathbb{E}\big{[}\sum_{t=1}^{T}u(x^{t},a^{t})\big{]}$ $\geq\underline{\mathrm{OBJ}}^{\mathcal{R}}\big{(}\frac{\mathrm{CReg}(T)}{T}\big{)}\geq U^{*}-\tfrac{4R}{\sqrt{G}}\sqrt{\frac{\mathrm{CReg}(T)}{T}}$ using a fixed contract in all rounds.
  • with a contextual no-swap-regret learning agent, the principal cannot obtain utility more than $\frac{1}{T}\mathbb{E}\big{[}\sum_{t=1}^{T}u(x^{t},a^{t})\big{]}\leq\overline{\mathrm{OBJ}}^{\mathcal{D}}\big{(}\frac{\mathrm{CSReg}(T)}{T}\big{)}\leq U^{*}+\tfrac{R}{G}\frac{\mathrm{CSReg}(T)}{T}$ even using time-varying contracts.
Providing the quantitative lower and upper bounds, the above results refine the result in Guruganesh et al. (2024) that the principal can obtain utility at least $U^{*}-o(1)$ against a no-regret learning agent and no more than $U^{*}+o(1)$ against a no-swap-regret agent. This again demonstrates the versatility of our general framework.
6 Discussion
In conclusion, our work showed that the principal in a generalized principal-agent problem with a no-regret learning agent can achieve the same and sometimes higher utility than the classic best-response model, while the agent’s no-swap-regret learning can cap the principal’s utility.
What is the intuition behind the difference between no-regret and no-swap-regret? If the principal only uses a fixed strategy through the interactions with the agent, then both no-regret and no-swap-regret algorithms ensure that the agent is approximately best responding to the principal’s fixed strategy; no-regret and no-swap-regret are not very different in this case. When the principal uses time-varying strategies, however, no-regret and no-swap-regret are different. We should think of the signal $s^{t}$ from the principal and the action $a^{t}$ recommended by the receiver’s algorithm together as a signal $(s^{t},a^{t})$ from some hypothetical signaling scheme $\pi^{\prime}$. No-swap-regret algorithm ensures that, whenever the signaling scheme $\pi^{\prime}$ sends $(s^{t},a^{t})$, it is almost optimal for the receiver to take the recommended action $a^{t}$. But no-regret only ensures that the agent will not regret when compared to taking any fixed action in all the rounds where signal $s^{t}$ is sent. The agent might regret if they could have taken different actions when receiving different $a^{t}$ signals (under the same $s^{t}$ signal). This means that a no-regret agent is not approximately-best-responding when the principal uses time-varying strategies, which explains why the principal can do better than the (approximately-)best-responding model. No-swap-regret agent, on the other hand, is always approximately-best-responding.
An interesting future direction is to characterize generalized principal-agent problems with a learning agent with private information. Our results only apply to complete-information settings. The conclusion that no-swap-regret learning can cap the principal’s utility is not true if the agent has private information, as shown by Mansour et al. (2022) in Bayesian Stackelberg games.
Appendix A Proof of Proposition 2.1
Let $\mathcal{A}$ be an arbitrary no-regret (no-swap-regret) learning algorithm for a multi-armed bandit (MAB) problem with $|A|$ arms. There exist such algorithms with regret $O(\sqrt{T|A|\log|A|})$ (variants of Exp3 (Auer et al., 2002)) and even $O(\sqrt{T|A|})$ (doubling trick + polyINF (Audibert and Bubeck, 2010)) for any time horizon $T>0$. By swap-to-external regret reductions, they can be converted to multi-armed bandit algorithms with swap regret $O(\sqrt{T|A|^{3}\log|A|})$ (Blum and Mansour, 2007) and $O(|A|\sqrt{T})$ (Ito, 2020). We then convert $\mathcal{A}$ into a contextual no-regret (contextual no-swap-regret) algorithm, in the following way:
Input: MAB algortihm $\mathcal{A}$. Arm set $A$. Context set $S$.
Instantiate $|S|$ copies $\mathcal{A}_{1},\ldots,\mathcal{A}_{|S|}$ of $\mathcal{A}$, and initialize their round number by $t_{1}=\cdots=t_{|S|}=0$.
for round $t=1,2,\ldots$ do
       Receive context $s^{t}$. Call $\mathcal{A}_{s^{t}}$ to obtain an action $a^{t}$.
       Play $a^{t}$ and obtain feedback (which includes the reward $v^{t}(a^{t})$ of action $a^{t}$).
       Feed the feedback to $\mathcal{A}_{s^{t}}$. Increase its round number $t_{s^{t}}$ by $1$.
end for
Algorithm 1 Convert any MAB algorithm to a contextual MAB algorithm
Proposition A.1.
The contextual regret of Algorithm 1 is at most
$\mathrm{CReg}(T)\leq\max\Big{\{}\sum_{s=1}^{|S|}\mathrm{Reg}(T_{s})~{}\Big{|}~{}T_{1}+\cdots+T_{|S|}=T\Big{\}},$
where $\mathrm{Reg}(T_{s})$ is the regret of $\mathcal{A}$ for time horizon $T_{s}$.
The contextual swap-regret of Algorithm 1 is at most
$\mathrm{CSReg}(T)\leq\max\Big{\{}\sum_{s=1}^{|S|}\mathrm{SReg}(T_{s})~{}\Big{|}~{}T_{1}+\cdots+T_{|S|}=T\Big{\}},$
where $\mathrm{SReg}(T_{s})$ is the swap-regret of $\mathcal{A}$ for time horizon $T_{s}$.
When plugging in $\mathrm{Reg}(T_{s})=O(\sqrt{|A|T_{s}})$, we obtain $\mathrm{CReg}(T)\leq O(\sqrt{|A||S|T})$.
When plugging in $\mathrm{SReg}(T_{s})=O(|A|\sqrt{T_{s}})$, we obtain $\mathrm{CSReg}(T)\leq O(|A|\sqrt{|S|T})$.
Proof.
The contextual regret of Algorithm 1 is
$\mathrm{CReg}(T)$ $=\max_{d:S\to A}\mathbb{E}\Big{[}\sum_{t=1}^{T}\big{(}v^{t}(d(s^{t}))-v^{t}(a^{t})\big{)}\Big{]}$
$=\max_{d:S\to A}\mathbb{E}\Big{[}\sum_{s=1}^{|S|}\sum_{t:s^{t}=s}\big{(}v^{t}(d(s))-v^{t}(a^{t})\big{)}\Big{]}$
$\leq\sum_{s=1}^{|S|}\max_{a^{\prime}\in A}\mathbb{E}\Big{[}\sum_{t:s^{t}=s}\big{(}v^{t}(a^{\prime})-v^{t}(a^{t})\big{)}\Big{]}$
$\leq\sum_{s=1}^{|S|}\mathbb{E}_{T_{s}}\big{[}\mathrm{Reg}(T_{s})\big{]}\quad\text{where $T_{s}$ is the number of rounds where $s^{t}=s$}$
$\leq\max\Big{\{}\sum_{s=1}^{|S|}\mathrm{Reg}(T_{s})~{}\Big{|}~{}T_{1}+\cdots+T_{|S|}=T\Big{\}}.$
When $\mathrm{Reg}(T_{s})=O(\sqrt{|A|T_{s}})$, by Jensen’s inequality we obtain
$\mathrm{CReg}(T)$ $\leq\sum_{s=1}^{|S|}O(\sqrt{|A|T_{s}})\leq O(\sqrt{|A|})\sqrt{|S|}\sqrt{\sum_{s=1}^{|S|}T_{s}}=O(\sqrt{|A||S|T}).$
The argument for contextual swap-regret is similar:
$\mathrm{CSReg}(T)$ $=\max_{d:S\times A\to A}\mathbb{E}\Big{[}\sum_{t=1}^{T}\big{(}v^{t}(d(s^{t},a^{t}))-v^{t}(a^{t})\big{)}\Big{]}$
$=\max_{d:S\times A\to A}\mathbb{E}\Big{[}\sum_{s=1}^{|S|}\sum_{t:s^{t}=s}\big{(}v^{t}(d(s,a^{t}))-v^{t}(a^{t})\big{)}\Big{]}$
$\leq\sum_{s=1}^{|S|}\max_{d^{\prime}:A\to A}\mathbb{E}\Big{[}\sum_{t:s^{t}=s}\big{(}v^{t}(d^{\prime}(a^{t}))-v^{t}(a^{t})\big{)}\Big{]}$
$\leq\sum_{s=1}^{|S|}\mathbb{E}_{T_{s}}\big{[}\mathrm{SReg}(T_{s})\big{]}\quad\text{where $T_{s}$ is the number of rounds where $s^{t}=s$}$
$\leq\max\Big{\{}\sum_{s=1}^{|S|}\mathrm{SReg}(T_{s})~{}\Big{|}~{}T_{1}+\cdots+T_{|S|}=T\Big{\}}.$
When $\mathrm{SReg}(T_{s})=O(|A|\sqrt{T_{s}})$, by Jensen’s inequality we obtain
$\mathrm{CSReg}(T)\leq\sum_{s=1}^{|S|}O(|A|\sqrt{T_{s}})\leq O(|A|)\sqrt{|S|}\sqrt{\sum_{s=1}^{|S|}T_{s}}=O(|A|\sqrt{|S|T}).$
Appendix B Proof of Example 3.1
Consider the quantal response model. Let $\gamma=\frac{\log(|A|\lambda)}{\lambda}$. Given signal $s$, with posterior $\mu_{s}$, we say an action $a\in A$ is not $\gamma$-optimal for posterior $\mu_{s}$ if
$v(\mu_{s},a^{*}_{s})-v(\mu_{s},a)\geq\gamma$
where $a^{*}_{s}$ is an optimal action for $\mu_{s}$. The probability that the receiver chooses not $\gamma$-optimal action $a$ is at most:
$\frac{\exp(\lambda v(\mu_{s},a))}{\sum_{a\in A}\exp(\lambda v(\mu_{s},a))}\leq\frac{\exp(\lambda v(\mu_{s},a))}{\exp(\lambda v(\mu_{s},a^{*}_{s}))}=\exp\Big{(}-\lambda\big{[}v(\mu_{s},a^{*}_{s})-v(\mu_{s},a)\big{]}\Big{)}\leq\exp(-\lambda\gamma)=\frac{1}{|A|\lambda}.$
By a union bound, the probability that the receiver chooses any not $\gamma$-approxiamtely optimal action is at most $\frac{1}{\lambda}$. So, the expected loss of utility of the receiver due to not taking the optimal action is at most
$(1-\frac{1}{\lambda})\cdot\gamma+\frac{1}{\lambda}\cdot 1\leq\frac{\log(|A|\lambda)+1}{\lambda}$
This means that the quantal response strategy is a $\frac{\log(|A|\lambda)+1}{\lambda}$-best-responding randomized strategy.
Consider inaccurate belief. Given signal $s$, the receiver has belief $\mu_{s}^{\prime}$ with total variation distance $d_{\mathrm{TV}}(\mu_{s}^{\prime},\mu_{s})\leq\varepsilon$ to the true posterior $\mu_{s}$. For any action $a\in A$, the difference of expected utility of action $a$ under beliefs $\mu_{s}^{\prime}$ and $\mu_{s}$ is at most $\varepsilon$:
$\big{|}\mathbb{E}_{\omega\sim\mu_{s}^{\prime}}[v(\omega,a)]-\mathbb{E}_{\omega\sim\mu_{s}}[v(\omega,a)]\big{|}\leq d_{\mathrm{TV}}(\mu_{s}^{\prime},\mu_{s})\leq\varepsilon.$
So, the optimal action for $\mu_{s}^{\prime}$ is a $2\varepsilon$-optimal action for $\mu_{s}$. This means that the receiver strategy is a deterministic $2\varepsilon$-best-responding strategy.
Appendix C Missing Proofs from Section 4
C.1 Proof of Claim 4.1
If no $G>0$ satisfies the claim, then there must exist an $a_{0}\in A$ such that for all $x\in\mathcal{X}$, $v(a_{0},\mu)-v(a^{\prime},\mu)\leq 0$ for some $a^{\prime}\in A\setminus\{a_{0}\}$. Namely,
$\max_{x\in\mathcal{X}}\min_{a^{\prime}\in A\setminus\{a_{0}\}}\big{\{}v(x,a_{0})-v(x,a^{\prime})\big{\}}\leq 0.$
Then, by the minimax theorem, we have
$\min_{\alpha^{\prime}\in\Delta(A\setminus\{a_{0}\})}\max_{x\in\mathcal{X}}\big{\{}v(x,a_{0})-v(x,\alpha^{\prime})\big{\}}=\max_{x\in\mathcal{X}}\min_{a^{\prime}\in A\setminus\{a_{0}\}}\big{\{}v(x,a_{0})-v(x,a^{\prime})\big{\}}\leq 0.$
This means that $a_{0}$ is weakly dominated by some mixed action $\alpha^{\prime}\in\Delta(A\setminus\{a_{0}\})$, violating Assumption 4.1.
C.2 Proof of Lemma 4.4
Let $A_{\Delta}(x)=\big{\{}a\in A\mid v(x,a)\geq v(x,a^{\prime})-\Delta,\forall a^{\prime}\in A\big{\}}$ be the set of $\Delta$-optimal actions of the agent in response to principal decision $x\in\mathcal{X}$. The proof of Lemma 4.4 uses another lemma that relates the principal utility under a randomized $\delta$-best-responding agent strategy $\rho\in\mathcal{R}_{\delta}(\pi)$ and that under an agent strategy $\rho^{\prime}$ that only randomizes over $A_{\Delta}(x_{s})$.
Lemma C.1.
Let $\pi=\{(\pi_{s},x_{s})\}_{s\in S}$ be a principal strategy and $\rho\in\mathcal{R}_{\delta}(\pi)$ be a randomized $\delta$-best-response to $\pi$. For any $\Delta>0$, there exists an agent strategy $\rho^{\prime}:s\mapsto\Delta(A_{\Delta}(x_{s}))$ that randomizes over $\Delta$-optimal actions only for each $x_{s}$, such that the principal’s utility under $\rho^{\prime}$ and $\rho$ satisfies: $\big{|}U(\pi,\rho^{\prime})-U(\pi,\rho)\big{|}\leq\frac{2B\delta}{\Delta}$.
Proof.
Let $a^{*}_{s}=\max_{a\in A}v(x_{s},a)$ be the agent’s optimal action for $x_{s}$. Let $\overline{A_{\Delta}(x_{s})}=A\setminus A_{\Delta}(x_{s})$ be the set of actions that are not $\Delta$-optimal for $x_{s}$. By the definition that $\rho\in\mathcal{R}_{\delta}(\pi)$ is a $\delta$-best-response to $\pi$, we have
$\quad\delta$ $\geq\sum_{s\in S}\pi_{s}\Big{[}v(x_{s},a^{*}_{s})-\sum_{a\in A}\rho(a|s)v(x_{s},a)\Big{]}$
$=\sum_{s\in S}\pi_{s}\Big{(}\sum_{a\in A_{\Delta}(x_{s})}\rho(a|s)\big{[}\underbrace{v(x_{s},a^{*}_{s})-v(x_{s},a)}_{\geq 0}\big{]}+\sum_{a\in\overline{A_{\Delta}(x_{s})}}\rho(a|s)\big{[}\underbrace{v(x_{s},a^{*}_{s})-v(x_{s},a)}_{>\Delta}\big{]}\Big{)}$
$\geq 0+\Delta\sum_{s\in S}\pi_{s}\sum_{a\in\overline{A_{\Delta}(x_{s})}}\rho(a|s)$
$=\Delta\sum_{s\in S}\pi_{s}\rho(\overline{A_{\Delta}(x_{s})}\,|\,s).$
Rearranging,
$\sum_{s\in S}\pi_{s}\rho(\overline{A_{\Delta}(x_{s})}\,|\,s)\leq\frac{\delta}{\Delta}.$ (35)
Then, we consider the randomized strategy $\rho^{\prime}$ that, for each $s$, chooses each action $a\in A_{\Delta}(x_{s})$ with the conditional probability that $\rho$ chooses $a$ given $a\in A_{\Delta}(x_{s})$:
$\rho^{\prime}(a\,|\,s)=\frac{\rho(a\,|\,s)}{\rho(A_{\Delta}(x_{s})\,|\,s)}.$
The sender’s utility under $\rho^{\prime}$ is:
$U(\pi,\rho^{\prime})=\sum_{s\in S}\pi_{s}\sum_{a\in A_{\Delta}(x_{s})}\frac{\rho(a\,|\,s)}{\rho(A_{\Delta}(x_{s})\,|\,s)}u(x_{s},a).$
The sender’s utility under $\rho$ is
$U(\pi,\rho)=\sum_{s\in S}\pi_{s}\sum_{a\in A_{\Delta}(x_{s})}\rho(a\,|\,s)u(x_{s},a)~{}+~{}\sum_{s\in S}\pi_{s}\sum_{a\in\overline{A_{\Delta}(x_{s})}}\rho(a\,|\,s)u(x_{s},a)$
Taking the difference between the two utilities, we get
$\big{|}U(\pi,\rho^{\prime})-U(\pi,\rho)\big{|}$
$\leq\Big{|}\sum_{s\in S}\pi_{s}\Big{(}\frac{1}{\rho(A_{\Delta}(x_{s})\,|\,s)}-1\Big{)}\sum_{a\in A_{\Delta}(x_{s})}\rho(a\,|\,s)u(x_{s},a)\Big{|}~{}+~{}\Big{|}\sum_{s\in S}\pi_{s}\sum_{a\in\overline{A_{\Delta}(x_{s})}}\rho(a\,|\,s)u(x_{s},a)\Big{|}$
$=\Big{|}\sum_{s\in S}\pi_{s}\frac{1-\rho(A_{\Delta}(x_{s})\,|\,s)}{\rho(A_{\Delta}(x_{s})\,|\,s)}\sum_{a\in A_{\Delta}(x_{s})}\rho(a\,|\,s)u(x_{s},a)\Big{|}~{}+~{}\Big{|}\sum_{s\in S}\pi_{s}\sum_{a\in\overline{A_{\Delta}(x_{s})}}\rho(a\,|\,s)u(x_{s},a)\Big{|}$
$\leq\sum_{s\in S}\pi_{s}\frac{1-\rho(A_{\Delta}(x_{s})\,|\,s)}{\rho(A_{\Delta}(x_{s})\,|\,s)}\sum_{a\in A_{\Delta}(x_{s})}\rho(a\,|\,s)\cdot B~{}+~{}\sum_{s\in S}\pi_{s}\sum_{a\in\overline{A_{\Delta}(\mu_{s})}}\rho(a\,|\,s)\cdot B$
$=B\sum_{s\in S}\pi_{s}\frac{\rho(\overline{A_{\Delta}(x_{s})}\,|\,s)}{\rho(A_{\Delta}(x_{s})\,|\,s)}\rho(A_{\Delta}(x_{s})\,|\,s)~{}+~{}B\sum_{s\in S}\pi_{s}\rho(\overline{A_{\Delta}(x_{s})}\,|\,s)$
[…]
This proves the lemma. ∎
We now prove Lemma 4.4.
Proof of Lemma 4.4.
Consider the objective $\underline{\mathrm{OBJ}}^{\mathcal{R}}(\delta)=\sup_{\pi}\min_{\rho\in\mathcal{R}_{\delta}(\pi)}U(\pi,\rho)$. By Lemma C.1, for any $(\pi,\rho)$ there exists an agent strategy $\rho^{\prime}:s\mapsto\Delta(A_{\Delta}(x_{s}))$ that only randomizes over $\Delta$-optimal actions such that $\big{|}U(\pi,\rho^{\prime})-U(\pi,\rho)\big{|}\leq\frac{2B\delta}{\Delta}$. Because minimizing over $\Delta(A_{\Delta}(x_{s}))$ is equivalent to minimizing over $A_{\Delta}(x_{s})$, which corresponds to deterministic $\Delta$-best-responding strategies, we get:
$\underline{\mathrm{OBJ}}^{\mathcal{R}}(\delta)\,=\,\sup_{\pi}\min_{\rho\in\mathcal{R}_{\delta}(\pi)}U(\pi,\rho)$ $\,\geq\,\sup_{\pi}\min_{\rho^{\prime}:s\mapsto\Delta(A_{\Delta}(x_{s}))}U(\pi,\rho^{\prime})-\tfrac{2B\delta}{\Delta}$
$\,=\,\sup_{\pi}\min_{\rho^{\prime}:s\mapsto A_{\Delta}(x_{s})}U(\pi,\rho^{\prime})-\tfrac{2B\delta}{\Delta}$
$\,=\,\underline{\mathrm{OBJ}}^{\mathcal{D}}(\Delta)-\tfrac{2B\delta}{\Delta}.$
C.3 Proof of Example 4.1
First, the sender’s optimal utility when the receiver exactly best responds is $2\mu_{0}$:
$U^{*}=2\mu_{0}.$
This is achieved by decomposing the prior $\mu_{0}$ into two posteriors $\mu_{a}=\frac{1}{2}$ and $\mu_{b}=0$ with probability $2\mu_{0}$ and $1-2\mu_{0}$ respectively, with the receiver taking action $a$ under posterior $\mu_{a}$ and $b$ under $\mu_{b}$.
Then, consider any signaling scheme of the sender, $\pi=\{(\pi_{s},\mu_{s})\}_{s\in S}$, which is a decomposition of the prior $\mu_{0}$ into $|S|$ posteriors $\mu_{s}\in[0,1]$ such that $\sum_{s\in S}\pi_{s}\mu_{s}=\mu_{0}$. Let $\rho:S\to\Delta(A)$ be a randomized strategy of the receiver, where $\rho(a|s)$ (and $\rho(b|s)$) denotes the probability that the receiver takes action $a$ (and $b$) under signal $s$. The sender’s expected utility under $\pi$ and $\rho$ is:
$U(\pi,\rho)=\sum_{s\in S}\pi_{s}\big{[}\rho(a|s)\cdot 1+\rho(b|s)\cdot 0\big{]}=\sum_{s\in S}\pi_{s}\rho(a|s).$ (36)
The receiver’s utility when taking action $a$ at posterior $\mu_{s}$ is $\mu_{s}\cdot 1+(1-\mu_{s})\cdot(-1)=2\mu_{s}-1$. So, the receiver’s expected utility under $\pi$ and $\rho$ is
$V(\pi,\rho)=\sum_{s\in S}\pi_{s}\big{[}\rho(a|s)\cdot(2\mu_{s}-1)+\rho(b|s)\cdot 0\big{]}=\sum_{s\in S}\pi_{s}\rho(a|s)(2\mu_{s}-1).$ (37)
Clearly, the receiver’s best response $\rho^{*}$ is to take action $a$ with certainty if and only if $\mu_{s}>\frac{1}{2}$, with expected utility
$V(\pi,\rho^{*})=\sum_{s:\mu_{s}>\frac{1}{2}}\pi_{s}(2\mu_{s}-1).$ (38)
To find $\underline{\mathrm{OBJ}}^{\mathcal{R}}(\delta)=\sup_{\pi}\min_{\rho\in\mathcal{R}_{\delta}(\pi)}U(\pi,\rho)$, we fix any $\pi$ and solve the inner optimization problem (minimizing the sender’s utility) regarding $\rho$:
$\min_{\rho}$ $U(\pi,\rho)=\sum_{s\in S}\pi_{s}\rho(a|s)$
s.t. $\rho\in\mathcal{R}_{\delta}(\pi)\quad\;\Longleftrightarrow\;\quad\delta$ $\geq V(\pi,\rho^{*})-V(\pi,\rho)$
$=\sum_{s:\mu_{s}>\frac{1}{2}}\pi_{s}(2\mu_{s}-1)-\sum_{s\in S}\pi_{s}\rho(a|s)(2\mu_{s}-1).$
Without loss of generality, we can assume that the solution $\rho$ satisfies $\rho(a|s)=0$ whenever $\mu_{s}\leq\frac{1}{2}$ (if $\rho(a|s)>0$ for some $\mu_{s}\leq\frac{1}{2}$, then making $\rho(a|s)$ to be 00 can decrease the objective $\sum_{s\in S}\pi_{s}\rho(a|s)$ while still satisfying the constraint). So, the optimization problem can be simplified to:
$\min_{\rho}$ $U(\pi,\rho)=\sum_{s:\mu_{s}>\frac{1}{2}}\pi_{s}\rho(a|s)$
s.t. $\delta\geq\sum_{s:\mu_{s}>\frac{1}{2}}\pi_{s}(2\mu_{s}-1)-\sum_{s:\mu_{s}>\frac{1}{2}}\pi_{s}\rho(a|s)(2\mu_{s}-1)$
$\quad\quad=\sum_{s:\mu_{s}>\frac{1}{2}}\pi_{s}(2\mu_{s}-1)(1-\rho(a|s)),$
$\rho(a|s)\in[0,1],~{}~{}~{}\forall s\in S:\mu_{s}>\tfrac{1}{2}.$
We note that this is a fractional knapsack linear program, which has a greedy solution (e.g., (Korte and Vygen, 2012)): sort the signals with $\mu_{s}>\frac{1}{2}$ in increasing order of $12\mu_{s}-$ 1 (equivalently, increasing order of $\mu_{s}$); label those signals by $s=1,\ldots,n$; find the first position $k$ for which $\sum_{s=1}^{k}\pi_{s}(2\mu_{s}-1)>\delta$:
$k=\min\big{\{}j:\sum_{s=1}^{j}\pi_{s}(2\mu_{s}-1)>\delta\big{\}};$
then, an optimal solution $\rho$ is given by:
$\begin{cases}\rho(a|s)=0&\text{for }s=1,\ldots,k-1;\\ \rho(a|k)=1-\frac{\delta-\sum_{s=1}^{k-1}\pi_{s}(2\mu_{s}-1)}{\pi_{k}(2\mu_{k}-1)}&\text{for }s=k;\\ \rho(a|s)=1&\text{for }s=k+1,\ldots,n.\end{cases}$
The objective value (sender’s expected utility) of the above solution $\rho$ is
$U(\pi,\rho)$ $=\sum_{s:\mu_{s}>\frac{1}{2}}\pi_{s}\rho(a|s)$
$=\pi_{k}\Big{(}1-\frac{\delta-\sum_{s=1}^{k-1}\pi_{s}(2\mu_{s}-1)}{\pi_{k}(2\mu_{k}-1)}\Big{)}+\sum_{s=k+1}^{n}\pi_{s}$
$=\sum_{s=k}^{n}\pi_{s}-\frac{\delta}{2\mu_{k}-1}+\sum_{s=1}^{k-1}\frac{\pi_{s}(2\mu_{s}-1)}{2\mu_{k}-1}.$
Since the signaling scheme $\pi$ must satisfy $\sum_{s\in S}\pi_{s}\mu_{s}=\mu_{0}$, we have
$\mu_{0}=\sum_{s\in S}\pi_{s}\mu_{s}\geq\sum_{s=1}^{n}\pi_{s}\mu_{s}=\sum_{s=1}^{k-1}\pi_{s}\mu_{s}+\sum_{s=k}^{n}\pi_{s}\mu_{s}\geq\sum_{s=1}^{k-1}\pi_{s}\mu_{s}+\sum_{s=k}^{n}\pi_{s}\mu_{k}$
$\implies\quad\sum_{s=k}^{n}\pi_{s}\leq\frac{\mu_{0}-\sum_{s=1}^{k-1}\pi_{s}\mu_{s}}{\mu_{k}}.$
So,
$U(\pi,\rho)$ $\leq\frac{\mu_{0}-\sum_{s=1}^{k-1}\pi_{s}\mu_{s}}{\mu_{k}}-\frac{\delta}{2\mu_{k}-1}+\sum_{s=1}^{k-1}\frac{\pi_{s}(2\mu_{s}-1)}{2\mu_{k}-1}$
$=\frac{\mu_{0}}{\mu_{k}}-\frac{\delta}{2\mu_{k}-1}+\sum_{s=1}^{k-1}\pi_{s}\Big{(}\frac{2\mu_{s}-1}{2\mu_{k}-1}-\frac{\mu_{s}}{\mu_{k}}\Big{)}.$
Since $\frac{2\mu_{s}-1}{2\mu_{k}-1}-\frac{\mu_{s}}{\mu_{k}}=\frac{\mu_{s}-\mu_{k}}{(2\mu_{s}-1)\mu_{k}}\leq 0$ for any $s\leq k-1$, we get
$U(\pi,\rho)$ $\leq\frac{\mu_{0}}{\mu_{k}}-\frac{\delta}{2\mu_{k}-1}=f(\mu_{k}).$
We find the maximal value of $f(\mu_{k})=\frac{\mu_{0}}{\mu_{k}}-\frac{\delta}{2\mu_{k}-1}$. Take its derivative:
$f^{\prime}(\mu_{k})$ $=-\frac{\mu_{0}}{\mu_{k}^{2}}+\frac{2\delta}{(2\mu_{k}-1)^{2}}=\frac{\big{[}(\sqrt{2\delta}+2\sqrt{\mu_{0}})\mu_{k}-\sqrt{\mu}_{0}\big{]}\cdot\big{[}(\sqrt{2\delta}-2\sqrt{\mu_{0}})\mu_{k}+\sqrt{\mu}_{0}\big{]}}{\mu_{k}^{2}(2\mu_{k}-1)^{2}},$
which has two roots $\frac{\sqrt{\mu}_{0}}{\sqrt{2\delta}+2\sqrt{\mu_{0}}}<\frac{1}{2}$ and $\frac{\sqrt{\mu}_{0}}{2\sqrt{\mu_{0}}-\sqrt{2\delta}}\in(\frac{1}{2},1)$ when $0<\delta<\frac{\mu_{0}}{2}$. So, $f(x)$ is increasing in $[\frac{1}{2},\frac{\sqrt{\mu}_{0}}{2\sqrt{\mu_{0}}-\sqrt{2\delta}})$ and decreasing in $(\frac{\sqrt{\mu}_{0}}{2\sqrt{\mu_{0}}-\sqrt{2\delta}},1]$. Since $\mu_{k}>\frac{1}{2}$, $f(\mu_{k})$ is maximized at $\mu_{k}=\frac{\sqrt{\mu}_{0}}{2\sqrt{\mu_{0}}-\sqrt{2\delta}}$. This implies
$U(\pi,\rho)$ $\leq f\big{(}\tfrac{\sqrt{\mu}_{0}}{2\sqrt{\mu_{0}}-\sqrt{2\delta}}\big{)}=\frac{\mu_{0}}{\sqrt{\mu}_{0}}(2\sqrt{\mu}_{0}-\sqrt{2\delta})-\frac{\delta}{2\frac{\sqrt{\mu}_{0}}{2\sqrt{\mu_{0}}-\sqrt{2\delta}}-1}=2\mu_{0}-2\sqrt{2\mu_{0}\delta}+\delta.$
This holds for any $\pi$. So, $\underline{\mathrm{OBJ}}^{\mathcal{R}}(\delta)=\sup_{\pi}\min_{\rho\in\mathcal{R}_{\delta}(\pi)}U(\pi,\rho)\leq U^{*}-2\sqrt{2\mu_{0}\delta}+\delta=U^{*}-\Omega(\sqrt{\delta})$.
C.4 Proof of Lemma 4.5
Let $(\pi,\rho)$ be a pair of principal strategy and agent strategy that achieves the optimal principal utility with an exactly-best-responding agent, namely, $U(\pi,\rho)=U^{*}$. Without loss of generality $\rho$ can be assumed to be deterministic, $\rho:S\to A$. The strategy $\pi$ consists of pairs $\{(\pi_{s},x_{s})\}_{s\in S}$ that satisfy
$\sum_{s\in S}\pi_{s}x_{s}=:\mu_{0}\in\mathcal{C},$ (39)
and the action $a=\rho(s)$ is optimal for the agent with respect to $x_{s}$. We will construct another principal strategy $\pi^{\prime}$ such that, even if the agent chooses the worst $\delta$-best-responding strategy to $\pi^{\prime}$, the principal can still obtain utility arbitrarily close to $U^{*}-\big{(}L\mathrm{diam}(\mathcal{X};\ell_{1})+2B\tfrac{\mathrm{diam}(\mathcal{X})}{\mathrm{dist}(\mathcal{C},\partial\mathcal{X})}\big{)}\tfrac{\delta}{G}$.
To construct $\pi^{\prime}$ we do the following: For each signal $s\in S$, with corresponding action $a=\rho(s)$, by Claim 4.1 there exists $y_{a}\in\mathcal{X}$ such that $v(y_{a},a)-v(y_{a},a^{\prime})\geq G$ for any $a^{\prime}\neq a$. Let $\theta=\frac{\delta}{G}+\varepsilon\in[0,1]$ for arbitrarily small $\varepsilon>0$, and let $\tilde{x}_{s}$ be the convex combination of $x_{s}$ and $y_{\rho(s)}$ with weights $1-\theta,\theta$:
$\tilde{x}_{s}=(1-\theta)x_{s}+\theta y_{\rho(s)}.$ (40)
We note that $a=\rho(s)$ is the agent’s optimal action for $\tilde{x}_{s}$ and moreover it is better than any other action $a^{\prime}\neq a$ by more than $\delta$:
$v(\tilde{x}_{s},a)-v(\tilde{x}_{s},a^{\prime})$ $=(1-\theta)\big{[}\underbrace{v(x_{s},a)~{}-~{}v(x_{s},a^{\prime})}_{\geq 0\text{ because $a=\rho(s)$ is optimal for $x_{s}$}}\big{]}~{}+~{}\theta\big{[}\underbrace{v(y_{a},a)-v(y_{a},a^{\prime})}_{\geq G\text{ by our choice of $y_{a}$}}\big{]}$
$\geq 0+\theta G>\tfrac{\delta}{G}G=\delta.$ (41)
Let $\mu^{\prime}$ be the convex combination of $\{\tilde{x}_{s}\}_{s\in S}$ with weights $\{\pi_{s}\}_{s\in S}$:
$\mu^{\prime}=\sum_{s\in S}\pi_{s}\tilde{x}_{s}.$ (42)
Note that $\mu^{\prime}$ might not satisfy the constraint $\mu^{\prime}\in\mathcal{C}$. So, we want to find another vector $z\in\mathcal{X}$ and a coefficient $\eta\in[0,1]$ such that
$(1-\eta)\mu^{\prime}+\eta z\in\mathcal{C}.$ (43)
(If $\mu^{\prime}$ already satisfies $\mu^{\prime}\in\mathcal{C}$, then let $\eta=0$.) To do this, we consider the ray starting from $\mu^{\prime}$ pointing towards $\mu_{0}$: $\{\mu^{\prime}+t(\mu_{0}-\mu^{\prime})\mid t\geq 0\}$. Let $z$ be the intersection of the ray with the boundary of $\mathcal{X}$:
$z=\mu^{\prime}+t^{*}(\mu_{0}-\mu^{\prime}),\quad\quad t^{*}=\operatorname*{arg\,max}\{t\geq 0\mid\mu^{\prime}+t(\mu_{0}-\mu^{\prime})\in\mathcal{X}\}.$
Then, rearranging $z=\mu^{\prime}+t^{*}(\mu_{0}-\mu^{\prime})$, we get
$\tfrac{1}{t^{*}}(z-\mu^{\prime})=\mu_{0}-\mu^{\prime}\quad\;\Longleftrightarrow\;\quad(1-\tfrac{1}{t^{*}})\mu^{\prime}+\tfrac{1}{t^{*}}z=\mu_{0}\in\mathcal{C},$
which satisfies (43) with $\eta=\frac{1}{t^{*}}$. We then give an upper bound on $\eta=\frac{1}{t*}$:
Claim C.2.
$\eta=\tfrac{1}{t^{*}}\leq\tfrac{\mathrm{diam}(\mathcal{X})}{\mathrm{dist}(\mathcal{C},\partial\mathcal{X})}\theta$.
Proof.
On the one hand,
$\|\mu_{0}-\mu^{\prime}\|$ $=\big{\|}\sum_{s\in S}\pi_{s}x_{s}-\sum_{s\in S}\pi_{s}\tilde{x}_{s}\big{\|}=\big{\|}\sum_{s\in S}\pi_{s}\theta(y_{\rho(s)}-x_{s})\big{\|}$
$\leq\theta\sum_{s\in S}\pi_{s}\big{\|}y_{\rho(s)}-x_{s}\big{\|}\leq\theta\sum_{s\in S}\pi_{s}\cdot\mathrm{diam}(\mathcal{X})=\theta\cdot\mathrm{diam}(\mathcal{X}).$
On the other hand, because $z-\mu^{\prime}$ and $\mu_{0}-\mu^{\prime}$ are in the same direction, we have
$\|z-\mu^{\prime}\|=\|z-\mu_{0}\|+\|\mu_{0}-\mu^{\prime}\|\geq\|z-\mu_{0}\|\geq\mathrm{dist}(\mathcal{C},\partial\mathcal{X})$
because $\mu_{0}$ is in $\mathcal{C}$ and $z$ is on the boundary of $\mathcal{X}$. Therefore, $\eta=\tfrac{1}{t^{*}}=\tfrac{\|\mu_{0}-\mu^{\prime}\|}{\|z-\mu^{\prime}\|}\leq\tfrac{\mathrm{diam}(\mathcal{X})}{\mathrm{dist}(\mathcal{C},\partial\mathcal{X})}\theta$. ∎
The convex combinations (43) (42) define a new principal strategy $\pi^{\prime}$ with $|S|+1$ signals, consisting of $\tilde{x}_{s}$ with probability $(1-\eta)\pi_{s}$ and $z$ with probability $\eta$, satisfying $\sum_{s\in S}(1-\eta)\pi_{s}\tilde{x}_{s}+\eta z=\mu_{0}\in\mathcal{C}$. Consider the agent’s worst (for the principal) $\delta$-best-responding strategies $\rho^{\prime}$ to $\pi^{\prime}$:
$\rho^{\prime}\in\operatorname*{arg\,min}_{\rho\in\mathcal{D}_{\delta}(\pi^{\prime})}U(\pi^{\prime},\rho).$
We note that $\rho^{\prime}(\tilde{x}_{s})$ must be equal to $\rho(s)$ for each $s\in S$. This is because $a=\rho(s)$ is strictly better than any other action $a^{\prime}\neq a$ by a margin of $\delta$ (41), so $a$ is the only $\delta$-optimal action for $\tilde{x}_{s}$.
Then, the principal’s expected utility under $\pi^{\prime}$ and $\rho^{\prime}$ is
$U(\pi^{\prime},\rho^{\prime})$ =(43),(42)${{=}}(1-\eta)\sum_{s\in S}\pi_{s}u(\tilde{x}_{s},\rho^{\prime}(\tilde{x}_{s}))~{}+~{}\eta u(z,\rho^{\prime}(z))$ ) )
$\geq(1-\eta)\sum_{s\in S}\pi_{s}u(\tilde{x}_{s},\rho(s))~{}-~{}\eta B$
$\geq(1-\eta)\sum_{s\in S}\pi_{s}\Big{(}u(x_{s},\rho(s))-L\underbrace{\|\tilde{x}_{s}-x_{s}\|}_{=\|\theta(y_{\rho(s)}-x_{s})\|\leq\theta\mathrm{diam}(\mathcal{X})}\Big{)}~{}-~{}\eta B$
$\geq(1-\eta)U(\pi,\rho)-L\theta\mathrm{diam}(\mathcal{X})-\eta B$
$\geq U(\pi,\rho)-L\theta\mathrm{diam}(\mathcal{X})-2\eta B$
(Claim C.2) $\geq U(\pi,\rho)-L\theta\mathrm{diam}(\mathcal{X})-2B\tfrac{\mathrm{diam}(\mathcal{X})}{\mathrm{dist}(\mathcal{C},\partial\mathcal{X})}\theta$
$=U(\pi,\rho)-\big{(}L\mathrm{diam}(\mathcal{X})+2B\tfrac{\mathrm{diam}(\mathcal{X})}{\mathrm{dist}(\mathcal{C},\partial\mathcal{X})}\big{)}(\tfrac{\delta}{G}+\varepsilon)$
$=U^{*}-\big{(}L\mathrm{diam}(\mathcal{X})+2B\tfrac{\mathrm{diam}(\mathcal{X})}{\mathrm{dist}(\mathcal{C},\partial\mathcal{X})}\big{)}\tfrac{\delta}{G}-O(\varepsilon).$
So, we conclude that
$\underline{\mathrm{OBJ}}^{\mathcal{D}}(\delta)\,=\,\sup_{\pi}\min_{\rho\in\mathcal{D}_{\delta}(\pi)}U(\pi,\rho)$ $\,\geq\,\min_{\rho\in\mathcal{D}_{\delta}(\pi^{\prime})}U(\pi^{\prime},\rho)$
$\,=\,U(\pi^{\prime},\rho^{\prime})\,\geq\,U^{*}-\big{(}L\mathrm{diam}(\mathcal{X})+2B\tfrac{\mathrm{diam}(\mathcal{X})}{\mathrm{dist}(\mathcal{C},\partial\mathcal{X})}\big{)}\tfrac{\delta}{G}-O(\varepsilon).$
Letting $\varepsilon\to 0$ finishes the proof for the case with the constraint $\sum_{s\in S}\pi_{s}x_{s}\in\mathcal{C}$.
The case without $\sum_{s\in S}\pi_{s}x_{s}\in\mathcal{C}$ is proved by letting $\eta=0$ in the above argument.
C.5 Proof of Claim 4.8
On the one hand,
$\|\mu_{0}-\mu^{\prime}\|$ $=\big{\|}\sum_{s\in S}\pi_{s}x_{s}-\sum_{s\in S}\sum_{a\in A}\pi_{s}\rho(a|s)\tilde{x}_{s,a}\big{\|}=\big{\|}\sum_{s\in S}\sum_{a\in A}\pi_{s}\rho(a|s)\theta_{s,a}(y_{a}-x_{s})\big{\|}$
[…]
On the other hand, because $z-\mu^{\prime}$ and $\mu_{0}-\mu^{\prime}$ are in the same direction, we have
$\|z-\mu^{\prime}\|=\|z-\mu_{0}\|+\|\mu_{0}-\mu^{\prime}\|\geq\|z-\mu_{0}\|\geq\mathrm{dist}(\mathcal{C},\partial\mathcal{X})$
because $\mu_{0}$ is in $\mathcal{C}$ and $z$ is on the boundary of $\mathcal{X}$. Therefore, $\eta=\tfrac{1}{t^{*}}=\tfrac{\|\mu_{0}-\mu^{\prime}\|}{\|z-\mu^{\prime}\|}\leq\tfrac{\mathrm{diam}(\mathcal{X})}{\mathrm{dist}(\mathcal{C},\partial\mathcal{X})}\tfrac{\delta}{G}$.